### v1

# Linear arboricity of graphs with large minimum degree

Abstract
In 1980, Akiyama, Exoo, and Harary conjectured that any graph $G$ can be decomposed into at most $\lceil(\Delta(G)+1)/2\rceil$ linear forests. We confirm the conjecture for sufficiently large graphs with large minimum degree. Precisely, we show that for any given $10<\varepsilon<$ 1, there exists $n_{0}\in\mathbb{N}$ for which the following statement holds: If $G$ is a graph on $n\geq n_{0}$ vertices of minimum degree at least $(1+\varepsilon)n/2$, then $G$ can be decomposed into at most $\lceil(\Delta(G)+1)/2\rceil$ linear forests.
Keywords: Linear forest; linear arboricity; robust expander; Hamilton decomposition
1 Introduction
Unless otherwise stated explicitly, graphs in this paper are simple and finite. A linear forest is a graph consisting of vertex-disjoint paths. The linear arboricity of a graph $G$, denoted $\operatorname{la}(G)$, is the smallest number of linear forests its edges can be partitioned into. This notion was introduced by Harary [15] in 1970 as one of the covering invariants of graphs, and has been studied quite extensively since then. The linear arboricity of any graph $G$ of maximum degree $\Delta$ is known to be at least $\lceil\Delta/2\rceil$ as the edges incident with a maximum degree vertex needs to be covered by at least $\lceil\Delta/2\rceil$ distinct linear forests. On the other hand, $\operatorname{la}(G)$ is conjectured to be at most $\lceil(\Delta+1)/2\rceil$. The conjecture, known as the Linear Arboricity Conjecture, was proposed by Akiyama, Exoo and Harary [1] in 1980.
Conjecture 1.1 (Linear Arboricity Conjecture).
Every graph $G$ satisfies $\operatorname{la}(G)\leq\lceil(\Delta(G)+1)/2\rceil$.
The conjecture would determine the linear arboricity exactly for graphs of odd maximum degree, as in that case the upper bound equals the lower bound. For graphs of even maximum degree it would imply that the linear arboricity must be one of only two possible values. However, determining the exact value among these two choices is NP-complete [19]. For regular graphs $G$, the conjecture implies $\operatorname{la}(G)=\lceil(\Delta(G)+1)/2\rceil$.
In terms of general bounds on $\operatorname{la}(G)$ for graphs $G$ with maximum degree $\Delta$, the first result of this type was obtained by Alon in 1988 [2], which states that $\operatorname{la}(G)\leq\Delta/2+O(\Delta\cdot\frac{\log\log\Delta}{\log\Delta})$. The error term was improved to $O(\Delta^{2/3}\log^{1/3}\Delta)$ by Alon and Spencer in 1992 [3]. Recently, the error term has been further improved to $O(\Delta^{0.661})$ by Ferber, Fox and Jain [10], and then to $O(\Delta^{1/2}\log^{4}\Delta)$ by Lang and Postle [17]. The Linear Arboricity Conjecture was also well studied for classes of graphs including degenerate graphs, graphs with small maximum degree, and planar graphs. For example, see [5, 7, 9, 13, 18, 23, 25, 26]. In this paper, we are interested in studying the conjecture for graphs of large minimum degree. For regular graphs, Glock, Kühn, and Osthus [12] in 2016 proved the following result.
Theorem 1.2 ([12]).
There exists $n_{0}\in\mathds{N}$ such that every $d$-regular graph $G$ on $n\geq n_{0}$ vertices with $d\geq\left\lfloor\frac{n-1}{2}\right\rfloor$ has a decomposition into $\left\lceil\frac{d+1}{2}\right\rceil$ linear forests.
By relaxing the regularity requirement in the theorem above but slightly strengthening the minimum degree of the graph, we obtain the result below.
Theorem 1.3.
For any given $10<\varepsilon<$ 1, there exists $n_{0}\in\mathds{N}$ for which the following statement holds. If $G$ is a graph on $n\geq n_{0}$ vertices with $\delta(G)\geq\frac{1}{2}(1+\varepsilon)n$, then $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$.
It is worth mentioning that in [11], Girão, Granet, Kühn, and Osthus proved that such graphs $G$ as described in Theorem 1.3 can be decomposed into at most $\max\{\frac{\operatorname{odd}(G)}{2},\frac{\Delta(G)}{2}\}+o(n)$ paths, and so $\operatorname{la}(G)\leq\max\{\frac{\operatorname{odd}(G)}{2},\frac{\Delta(G)}{2}\}+o(n)$, where $\operatorname{odd}(G)$ is the number of odd degree vertices of $G$.
Our proof of Theorem 1.3 relies on the intriguing results of Kühn and Osthus [16] that regular robust expanders of linear even degree has a Hamilton decomposition, and of Csaba, Kühn, Lo, Osthus, and Treglown [6] that large $n$-vertex $d$-regular graphs with $d$ even and $d\geq\lfloor n/2\rfloor$ has a Hamilton decomposition. The proof of Theorem 1.3 will be given in Section 3, after we introduce some preliminaries in the next section.
2 Notation and Preliminaries
Let $G$ be a graph and $S\subseteq V(G)$. We let $e(G)=|E(G)|$. The subgraph of $G$ induced on $S$ and $G-S$ is respectively denoted by $G[S]$ and $G-S$. For $v\in V(G)$, we write $G-v$ for $G-\{v\}$. For two disjoint subsets $S,T$ of $V(G)$, we denote by $E_{G}(S,T)$ the set of edges in $G$ with one endvertex in $S$ and the other in $T$, and let $e_{G}(S,T)=|E_{G}(S,T)|$. If $F\subseteq E(G)$, then $G-F$ is obtained from $G$ by deleting all edges in $F$. If $F$ is a subset of edges from the complement $\overline{G}$, then $G+F$ is obtained from $G$ by adding the edges of $F$ to $G$. For two integers $p$ and $q$, let $[p,q]=\{i\in\mathbb{Z}:p\leq i\leq q\}$.
2.1 Degree sequences and number of maximum degree vertices
Let $n\geq 2$ be an integer. A sequence of nonnegative and non-increasing integers $(d_{1},\ldots,d_{n})$ is graphic if there is a graph $G$ such that the degree sequence of $G$ is $(d_{1},\ldots,d_{n})$. In this case, we say that $G$ realizes $(d_{1},\ldots,d_{n})$.
The first result below by Hakimi [14] gives a sufficient and necessary condition for a sequence of integers to be the degree sequence of a multigraph. The second result by the second author [21] will be used to construct a supergraph of a given graph $G$ by adding additional vertices.
Theorem 2.1 ([14]).
Let $0\leq d_{n}\leq\ldots\leq d_{1}$ be integers. Then there exists a multigraph $G$ on vertices $v_{1},v_{2},\ldots,v_{n}$ such that $d_{G}(v_{i})=d_{i}$ for all $i\in[1,n]$ if and only if $\sum\limits_{i=1}^{n}d_{i}$ is even and $\sum\limits_{i=2}^{n}d_{i}\geq d_{1}$.
Lemma 2.2 ([21]).
Let $n$ and $d$ be positive integers such that $n\geq d+1\geq 3$. Suppose $(d_{1},\ldots,d_{n})$ is a sequence of positive integers with $d_{1}=\ldots=d_{t}=d$ and $d_{t+1}=\ldots=d_{n}=d-1$ and $\sum\limits_{i=1}^{n}d_{i}$ even, where $t\in[1,n]$. Then $(d_{1},\ldots,d_{n})$ is graphic. Furthermore, the sequence can be realized in polynomial time in $n$.
We will also need the following result by Plantholt and the second author [20] which gives a lower bound on the number of maximum degree vertices if we assume that all vertices of degree less than maximum degree form a clique.
Lemma 2.3 ([20]).
Let $G$ be a graph on $n$ vertices such that all vertices of degree less than $\Delta(G)$ are mutually adjacent in $G$. Then $G$ has more than $n/2$ vertices of maximum degree.
2.2 Hamiltonicity and decomposition results on dense graphs
The first classic result below by Dirac from 1952 will be used to find a Hamilton path connecting two given vertices, and the second result by Plantholt and the second author [20] is to find a linear forest in $G$ with the endvertices prescribed as long as the size of these prescribed vertices is not too big.
Theorem 2.4 ([8]).
Let $G$ be a graph on $n\geq 3$ vertices. If $\delta(G)\geq n/2$, then $G$ has a Hamilton cycle; and if $\delta(G)\geq(n+1)/2$, then $G$ is Hamilton-connected.
Lemma 2.5 ([20]).
Let $10<1/n_{0}\ll\varepsilon<$ 1 and $G$ be a graph on $n\geq n_{0}$ vertices such that $\delta(G)\geq(1+\varepsilon)n/2$. Moreover, let $M=\{a_{1}b_{1},\ldots,a_{t}b_{t}\}$ be a matching in the complete graph on $V(G)$ of size at most $\varepsilon n/8$. Then there exist vertex-disjoint paths $P_{1},P_{2},\ldots,P_{t}$ in $G$ such that $\bigcup_{i=1}^{t}V(P_{i})=V(G)$, $P_{i}$ joins $a_{i}$ to $b_{i}$, and these paths can be found in polynomial time.
The following result by Csaba, Kühn, Lo, Osthus, and Treglown [6] is one of the most significant results established in the recent years, which implies the famous 1-Factorization Conjecture for large graphs, and was also used by Glock, Kühn and Osthus [12] to prove Theorem 1.2.
Theorem 2.6 ([6]).
There exists $n_{0}\in\mathds{N}$ such that whenever $n\geq n_{0},d\geq\left\lfloor\frac{n}{2}\right\rfloor$ and $G$ is a $d$-regular graph on $n$ vertices, then $G$ has a decomposition into Hamilton cycles and at most one perfect matching.
2.3 Matchings in almost-regular graphs
Let $G$ be a graph. A matching $M$ in $G$ is a set of vertex-disjoint edges. The set of vertices covered or saturated by $M$ is denoted by $V(M)$. We say that $M$ is perfect if $V(M)=V(G)$. The graph $G$ is factor-critical if $G$ does not have a perfect matching but $G-v$ has one for any $v\in V(G)$.
For a subset $X\subseteq V(G)$, we denote by $o(G-X)$ the number of components of $G-X$ that each have odd order. Define ${\rm df}(G)=\max_{X\subseteq V(G)}(o(G-X)-|X|)$. By applying Tutte’s Theorem [22] to the graph obtained from $G$ by adding ${\rm df}(G)$ vertices and joining all edges between $V(G)$ and these new vertices, Berge [4] observed that the maximum size of a matching in $G$ is $\frac{1}{2}(n(G)-{\rm df}(G))$. West [24] gave a short proof of this result. To state his result, we need some more definitions.
For $X\subseteq V(G)$, denote by $Y$ the set of components of $G-X$. We define an auxiliary bipartite multigraph $B(X)$ by contracting each component of $G-X$ to a single vertex and deleting edges within $X$. In other words, the multigraph $B(X)$ has an edge $xy$ for $x\in X$ and $y\in Y$ if and only if $x$ has a neighbor in $G$ in the component of $G-X$ corresponding to $y$. West proved Statements (1)-(2) below in [24]. Statement (3) follows easily by combining Statement (2) and two facts about matchings in bipartite graphs: Let $G[X,Y]$ be a bipartite graph, and let $d=\max\{d_{G}(x):x\in X\}$. First, $G$ has a matching saturating all vertices from $Y$ of degree at least $d$ in $G$ (by applying Hall’s Theorem to $G[X,Y_{d}]$, where $Y_{d}=\{y\in Y:d_{G}(y)\geq d\}$). Second, if $G$ has a matching $M_{1}$ saturating $X_{1}\subseteq X$ and a matching $M_{2}$ saturating $Y_{1}\subseteq Y$, then $G$ has a matching saturating $X_{1}\cup Y_{1}$ (by deleting some edges incident with vertices of degree 2 in the subgraph of $G$ induced on $M_{1}\cup M_{2}$ appropriately).
Lemma 2.7.
Let $G$ be a multigraph and $X\subseteq V(G)$ be maximal with $o(G-X)-|X|={\rm df}(G)$. Then the following statements hold.
  1. (1)
    ([24, Lemma 2]) Every component of $G-X$ has odd order, and the component is factor-critical.
  2. (2)
    ([24, Lemma 3]) The bipartite multigraph $B(X)$ has a matching covering $X$.
  3. (3)
    The bipartite multigraph $B(X)$ has a matching covering $X$ and all vertices of degree at least $d$ from $V(B(X))\setminus X$ in $B(X)$, where $d=\max\{d_{B(X)}(x):x\in X\}$.
A graph $G$ is almost $r$-regular for some integer $r\geq 0$ if except one vertex, all other vertices have degree $r$ in $G$, and the exceptional vertex has degree $r+1$ in $G$.
Lemma 2.8.
Let $G$ be an $n$-vertex almost $r$-regular graph for some integer $r\geq 0$. Then $\overline{G}$ has a matching covering at least $n-\frac{n}{n-r}-3$ vertices of $\overline{G}$ excluding the vertex of degree $r+1$ in $G$.
Proof.
Let $x\in V(G)$ be the vertex with $d_{G}(x)=r+1$. Note that $\overline{G}$ is almost $(n-1-r)$-regular. If $\overline{G}$ has a perfect matching, then we are done. Thus we suppose that $\overline{G}$ does not have a perfect matching. Then by Lemma 2.7, we choose $S\subseteq V(\overline{G})$ maximal such that $o(\overline{G}-S)-|S|={\rm df}(\overline{G})$. Then $\overline{G}-S$ has no even component by Lemma 2.7(1). For any odd component $D$ of $\overline{G}-S$ such that $x\not\in V(D)$, suppose that $|V(D)|\leq n-r-1$. Then we have
$e_{\overline{G}}(V(D),S)$$\geq$ $|V(D)|(n-r-1-(|V(D)|-1))$
$\geq$ $n-r-1,$
where the last inequality is obtained as the quadratic function $|V(D)|(n-r-1-(|V(D)|-1))$ in $|V(D)|$ achieves its minimum at $|V(D)|=1$ or $|V(D)|=n-r-1$. There are at most $\frac{n}{n-r}$ odd components $D$ of $\overline{G}-S$ with $|V(D)|\geq n-r$, and so there are at most $\frac{n}{n-r}+1$ odd components $D$ of $\overline{G}-S$ such that $e_{\overline{G}}(V(D),S)\leq n-r-2$. As every vertex of $S$ has degree at most $n-1-r$ in $\overline{G}$, by Lemma 2.7(3), we know that $o(\overline{G}-S)-|S|\leq\frac{n}{n-r}+1$ and so ${\rm df}(\overline{G})\leq\frac{n}{n-r}+1$. Thus $\overline{G}$ has a matching that covers at least $n-\frac{n}{n-r}-1$ vertices of $\overline{G}$, and so it has a matching that covers at least $n-\frac{n}{n-r}-3$ vertices of $\overline{G}$ excluding $x$. ∎
2.4 Robust expanders
Given $10<\nu\leq\tau<$ 1, we say that a graph $G$ on $n$ vertices is a robust $(\nu,\tau)$-expander, if for all $S\subseteq V(G)$ with $\tau n\leq|S|\leq(1-\tau)n$ the number of vertices that have at least $\nu n$ neighbors in $S$ is at least $|S|+\nu n$. The $\nu$-robust neighborhood $RN_{G}(S)$ is the set of all those vertices of $G$ which have at least $\nu n$ neighbors in $S$. The theorem below by Kühn and Osthus [16] is profound and also is a powerful tool that has been used to establish other results.
Theorem 2.9 ([16]).
For every $\alpha>0$ there exists $\tau=\tau(\alpha)>0$ such that for every $\nu>0$ there exists an integer $N_{0}=N_{0}(\alpha,\nu,\tau)$ for which the following holds.
  1. (i)
    $G$ is an $r$-regular graph on $n\geq N_{0}$ vertices, where $r\geq\alpha n$ is even;
  2. (ii)
    $G$ is a robust $(\nu,\tau)$-expander.
Then $G$ has a Hamilton decomposition. Moreover, this decomposition can be found in polynomial time in $n$.
Kühn and Osthus [16] also showed that graphs with large minimum degree are robust expanders.
Lemma 2.10 ([16]).
Suppose that $10<\nu\leq\tau\leq\varepsilon<$ 1 are such that $\varepsilon\geq 4\nu/\tau$. Let $G$ be a graph on $n$ vertices with minimum degree $\delta(G)\geq(1+\varepsilon)n/2$. Then $G$ is a robust $(\nu,\tau)$-expander.
3 Proof of Theorem 1.3
Proof of Theorem 1.3.
Let $\tau^{*}=\tau(1/4)$ and $N^{*}_{0}=N_{0}(1/4,\eta^{2},\tau^{*})$ be defined as in Theorem 2.9. For any given $\varepsilon$, we let
$0<1/n_{0}\leq 1/N_{0}^{*}\ll\eta\ll\tau^{*},\varepsilon,$ (1)
where $n_{0}$ is additionally taken to be at least the lower bound on the number of vertices required in Theorems 1.2 and 2.6. Let $G$ be a graph on $n\geq n_{0}$ vertices with $\delta(G)\geq(1+\varepsilon)n/2$, and let
$\Delta=\Delta(G),\quad\delta=\delta(G),\quad\text{and}\quad V_{i}(G)=\{v\in V(G):d_{G}(v)=i\}.$
We simply let $V_{\delta}=V_{\delta}(G),V_{\delta+1}=V_{\delta+1}(G)$, and $V_{\Delta}=V_{\Delta}(G)$.
For any two vertices of degree less than $\Delta$, if they are not adjacent in $G$, we add an edge joining them to $G$. By iterating this process, we thus assume that all vertices with degrees less than $\Delta$ are mutually adjacent in $G$. Thus by Lemma 2.3, we have
$|V_{\Delta}|>\frac{n}{2}.$ (2)
If $G$ is regular, then we are done by Theorem 1.2. Thus we assume that $G$ is not regular. Let
$g=\Delta-\delta,\quad U=\{v\in V(G):\Delta-d_{G}(v)\geq\eta n\},\quad W=V(G)\setminus(V_{\Delta}\cup V_{\delta}),$
and let $V(G)=\{v_{1},\ldots,v_{n}\}$. Assume, without loss of generality, that $d_{G}(v_{1})\leq\ldots\leq d_{G}(v_{n})$. We call a vertex from $W$ a middle-degree vertex. We make three general claims below that together will handle most of the cases in the proof.
Claim 3.1.
If $|V_{\delta}\cup V_{\delta+1}|<\eta n$ and $g\geq 3$, then we may assume that $|W|\leq 1$.
Proof.
Suppose to the contrary that $|W|\geq 2$. Let $u,v\in W$ be two distinct vertices, and $G^{*}=G-(V_{\delta}\cup(V_{\delta+1}\setminus\{u,v\}))$. As $|V_{\delta}\cup V_{\delta+1}|<\eta n$, $G^{*}$ is Hamilton-connected by Theorem 2.4. Let $P$ be a Hamilton $(u,v)$-path of $G^{*}$ and we consider $G-E(P)$. We have $\delta(G-E(P))=\delta$ and $\Delta(G-E(P))=\Delta-2\geq\delta+1$. Thus $\Delta(G-E(P))-\delta(G-E(P))\leq g-1$. Furthermore, $\operatorname{la}(G-E(P))\leq\lceil(\Delta-2+1)/2\rceil$ implies that $\operatorname{la}(G)\leq 1+\operatorname{la}(G-E(P))\leq\lceil(\Delta+1)/2\rceil$. By iterating this process of removing paths with endvertices of middle-degree but covering all vertices of degree at least $\delta+2$, we can reduce $G$ to a spanning subgraph $G^{\prime}$ of $G$ with $\delta(G^{\prime})=\delta$ such that either $|V_{\delta}(G^{\prime})\cup V_{\delta+1}(G^{\prime})|\geq\eta n$, $\Delta(G^{\prime})-\delta(G^{\prime})\leq 2$ or $G^{\prime}$ has at most one middle-degree vertex. As $G^{\prime}$ satisfies all the assumptions of Theorem 1.3, we can then consider $G^{\prime}$ in the place of $G$. We may thus assume that $G^{\prime}$ has at most one middle-degree vertex under the assumption that $|V_{\delta}(G^{\prime})\cup V_{\delta+1}(G^{\prime})|<\eta n$ and $\Delta(G^{\prime})-\delta(G^{\prime})\geq 3$. ∎
For a positive integer $d$, the $d$-deficiency of $v$ in $G$ is $\operatorname{def}_{G}(v,d)=\max\{d-d_{G}(v),0\}$.
Claim 3.2.
For some positive integer $d$, suppose that $\operatorname{def}_{G}(v_{1},d)\geq\ldots\geq\operatorname{def}_{G}(v_{n},d)$ and $\operatorname{def}_{G}(v_{1},d)\leq\eta n+1$. If $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},d)$ is even and $\operatorname{def}_{G}(v_{1},d)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},d)$, then there exists some integer $k\leq\varepsilon n/8$ such that after removing $k$ edge-disjoint linear forests from $G$, we can get a graph $G^{*}$ such that $d_{G^{*}}(v_{i})=d_{G}(v_{i})-2k+\operatorname{def}_{G}(v_{i},d)$ for each $i\in[1,n]$.
Proof.
By Theorem 2.1, there exists a multigraph $H$ on $V(G)$ such that $d_{H}(x_{i})=\operatorname{def}_{G}(v_{i},d)$ for each $i\in[1,n]$. The multigraph $H$ will aid us to find a desired subgraph $G^{*}$ of $G$. Note that $\Delta(H)=\operatorname{def}_{G}(v_{1},d)=d-d_{G}(v_{1})\leq\eta n+1$ and so $e(H)\leq(\eta n^{2}+n)/2$.
By Vizing’s Theorem on chromatic index, we have $\chi^{\prime}(H)\leq\Delta(H)+\mu(H)\leq 2\Delta(H)\leq 2\eta n+2$, where $\mu(H)$ is the maximum number of edges joining two vertices in $H$. Thus we can greedily partition the edges of $H$ into
$k$$\leq$ $e(H)/(\varepsilon n/16)\leq\frac{(\eta n^{2}+n)/2}{\varepsilon n/16}<\varepsilon n/8$
matchings $M_{1},M_{2},\ldots,M_{k}$ each of size at most $\varepsilon n/16$.
Now we take out linear forests from $G$ by applying Lemma 2.5 with $M_{1},M_{2},\ldots,M_{k}$. More precisely, we define spanning subgraphs $G_{0},G_{1},\ldots,G_{k}$ of $G$ and edge-disjoint linear forests $F_{1},F_{2},\ldots,F_{k}$ such that
  1. (a)
    $G_{0}:=G$ and $G_{i}:=G_{i-1}-E(F_{i})$ for $i\geq 1$.
  2. (b)
    $F_{i}$ is a spanning linear forest in $G_{i-1}$ whose leaves are precisely the vertices in $V(M_{i})$.
Let $G_{0}=G$ and suppose that for some $i\in[1,k]$, we have already defined $G_{0},G_{1},\ldots,G_{i-1}$ and $F_{1},\ldots,F_{i-1}$. Since $\Delta(F_{1}\cup\ldots\cup F_{i-1})\leq 2(i-1)\leq 2(k-1)\leq\varepsilon n/4$, it follows that $\delta(G_{i-1})\geq(1+\varepsilon)n/2-\varepsilon n/4\geq(1+0.5\varepsilon)n/2$. Since $M_{i}$ has size at most $\varepsilon n/16$, we can apply Lemma 2.5 to $G_{i-1}$ and $M_{i}$ to obtain a spanning linear forest $F_{i}$ in $G_{i-1}$ whose leaves are precisely the vertices in $V(M_{i})$. Set $G_{i}:=G_{i-1}-E(F_{i})$.
We claim that $G_{k}$ satisfies the degree constraint. Consider any vertex $u\in V(G_{k})$. For every $i\in[1,k]$, $d_{F_{i}}(u)=1$ if $u$ is an endvertex of some edge of $M_{i}$ and $d_{F_{i}}(u)=2$ otherwise. Since $M_{1},M_{2},\ldots,M_{k}$ partition $E(H)$, we know that $\sum\limits_{i=1}^{k}d_{F_{i}}(u)=2k-d_{H}(u)=2k-\operatorname{def}_{G}(u,d)$. Thus
$d_{G_{k}}(u)=d_{G}(u)-\sum\limits_{i=1}^{k}d_{F_{i}}(u)=d_{G}(u)-2k+\operatorname{def}_{G}(u,d).$
Thus letting $G^{*}=G_{k}$ gives us the desired subgraph. ∎
Claim 3.3.
Suppose that $|V_{\delta}\cup V_{\delta+1}|<\eta n$ and $|W|\leq 1$. Then $\operatorname{la}(G)\leq\lceil\frac{\Delta+1}{2}\rceil$.
Proof.
As $g=\Delta-\delta\leq n-1-(1+\varepsilon)n/2<n/2$, we have $|V_{\Delta}|>g$ by Inequality (2). If $|W|=1$, we let $w\in W$ and $g_{0}=d_{G}(w)-\delta$. We will remove from $G$ some edge-disjoint paths and reduce it to a graph whose maximum degree and minimum degree differ by exactly one. Recall that by our definition of $W$, we have $V_{\Delta}=V(G)\setminus(V_{\delta}\cup W)$.
We will select $2\lceil g/2\rceil$ vertices $x_{1},y_{1},\ldots,x_{\lceil g/2\rceil},y_{\lceil g/2\rceil}$ and take a Hamilton path between $x_{i}$ and $y_{i}$ for each $i\in[1,\lceil g/2\rceil]$, where we will let $w$ be exactly one of those selected vertices if $|W|=1$. We describe below how to select those vertices.
As $|V_{\delta}\cup V_{\delta+1}|<\eta n$ and $|W|\leq 1$, $G-V_{\delta}$ is Hamilton-connected by Theorem 2.4. If $|W|=1$ and $g_{0}$ is odd, we let $x_{1}=w$ and $y_{1}$ be any vertex from $V_{\Delta}$, and let $P_{1}$ be a Hamilton $(x_{1},y_{1})$-path of $G-V_{\delta}$. Otherwise, we let $x_{1},y_{1}\in V_{\Delta}$ be any two distinct vertices, and let $P_{1}$ be a Hamilton $(x_{1},y_{1})$-path of $G-V_{\delta}-W$.
Suppose $P_{1}=u_{1}u_{2}\ldots u_{p}$, where $p=n-|V_{\delta}|$ if $g_{0}$ is odd and $p=n-|V_{\delta}|-|W|$ otherwise, $u_{1}=x_{1}$ and $u_{p}=y_{1}$. Note that $p>g+1$. Now for each $i\in[2,\lceil g/2\rceil]$, we let $x_{i}=u_{2(i-1)}$ and $y_{i}=u_{2(i-1)+1}$. Now,
  • if $g_{0}\leq 1$, for each $i\in[2,\lceil g/2\rceil]$, we let $P_{i}$ be a Hamilton $(x_{i},y_{i})$-path of $G-V_{\delta}-W-\bigcup_{j=1}^{i-1}E(P_{j})$;
  • if $g_{0}\geq 2$, for each $i\in[2,\lceil(g_{0}+1)/2\rceil]$, we let $P_{i}$ be a Hamilton $(x_{i},y_{i})$-path of $G-V_{\delta}-\bigcup_{j=1}^{i-1}E(P_{j})$; for each $i\in[\lceil(g_{0}+1)/2\rceil+1,\lceil g/2\rceil]$, we let $P_{i}$ be a Hamilton $(x_{i},y_{i})$-path of $G-V_{\delta}-W-\bigcup_{j=1}^{i-1}E(P_{j})$.
The paths described above all exist as when $g_{0}\geq 2$, for each $i\in[2,\lceil(g_{0}+1)/2\rceil]$, we have $\delta(G-V_{\delta}-\bigcup_{j=1}^{i-1}E(P_{j}))\geq(1+\varepsilon)n/2+g_{0}-|V_{\delta}|-g_{0}>n/2$. Similarly, for each $i\in[\lceil(g_{0}+1)/2\rceil+1,\lceil g/2\rceil]$, we have $\delta(G-V_{\delta}-W-\bigcup_{j=1}^{i-1}E(P_{j}))\geq(1+\varepsilon)n/2+g-|V_{\delta}|-1-g>n/2$.
Let $G_{1}$ be the resulting graph, and $Z=\{x_{1},y_{1},\ldots,x_{\lceil g/2\rceil},y_{\lceil g/2\rceil}\}$. By the construction above, vertices of $V_{\delta}$ were not involved in the process. Thus we have $d_{G_{1}}(v)=\delta$ for each $v\in V_{\delta}$. Every vertex $v$ of $V_{\Delta}\setminus Z$ was used as an internal vertex for each of the path removed, thus we have $d_{G_{1}}(v)=\Delta-2\lceil g/2\rceil$. If $|W|=0$, then every vertex of $Z$ was used as an endvertex of exactly one of the paths removed, and as an internal vertex of all other paths, thus we have $d_{G_{1}}(v)=\Delta-2\lceil g/2\rceil+1$ for each $v\in Z$. If $|W|=1$, then the vertex $w$ was used exactly $\lceil g_{0}/2\rceil$ times, where exactly $\lceil(g_{0}+1)/2\rceil-1$ times as an internal vertex of $P_{2},\ldots,P_{\lceil(g_{0}+1)/2\rceil}$, and $\lceil g_{0}/2\rceil+1-\lceil(g_{0}+1)/2\rceil$ time as an endvertex of $P_{1}$ (when $g_{0}$ is odd). As a consequence, we have $d_{G_{1}}(w)=d_{G}(w)-2\lceil g_{0}/2\rceil+\lceil g_{0}/2\rceil+1-\lceil(g_{0}+1)/2\rceil=\delta$.
As edges of $E(P_{1})$ were removed from $G$ when getting $G_{1}$, we know that $E(P_{1})\subseteq E(\overline{G}_{1})$. By our choice of the vertices in $Z$, we know that $\overline{G}_{1}$ has a matching with edges in $u_{2\lceil g/2\rceil}P_{1}u_{p-1}$ covering at least $n-|V_{\delta}|-|Z|-1$ vertices of $G_{1}$ excluding vertices of $Z\cup V_{\delta}\cup W$, as all vertices from $V(P_{1})\setminus\{u_{1}\}$ are from $V_{\Delta}$. We consider two cases below regarding the parity of $g$ to complete the proof.
Case 1: $g$ is even.
Then we have $d_{G_{1}}(v)=\delta+1$ for every $v\in Z$ and $d_{G_{1}}(v)=\delta$ for every $v\in V(G)\setminus Z$. Thus $\Delta(G_{1})=\delta+1$. Since $g$ is even, we know that $\delta+1$ and $\Delta$ have different parities. .
If $\Delta$ is odd, then $\delta+1$ is even. As $G_{1}$ has precisely $n-g$ vertices of degree $\delta$, it follows that $n-g$ is even and so $n$ is even. Also, we know that $n-g\geq\Delta+1-g=\delta+1$. As $\overline{G}_{1}$ has a matching covering at least $n-|V_{\delta}|-|Z|-1$ vertices of $G_{1}$ excluding $Z$, it has a matching $M$ covering exactly $n-g-(\delta+1)$ vertices not including any vertex of $Z$. We let $G_{2}$ be obtained from $G_{1}+M$ by adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup V(M))$. The graph $G_{2}$ is $(\delta+1)$-regular and so has a Hamilton decomposition by Theorem 2.6. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+1)/2$ linear forests. Thus $\operatorname{la}(G)\leq g/2+(\delta+1)/2=(\Delta+1)/2$, as desired.
If $\Delta$ is even, then $\delta+1$ is odd. If $n$ is odd, then $n-g-(\delta+1)$ is even. We proceed exactly the same way above in getting a $(\delta+1)$-regular graph $G_{2}$. By Theorem 1.2, we have $\operatorname{la}(G_{2})\leq(\delta+2)/2$. Thus $\operatorname{la}(G)\leq g/2+(\delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired. Thus we assume that $n$ is even. As $n-g-\delta-2+n-\delta-2<n-g-\varepsilon n-4<n-|V_{\delta}|-|Z|-1$, $\overline{G}_{1}$ has vertex-disjoint matchings $M_{1}$ covering exactly $n-g-\delta-2$ vertices of $G_{1}$ and $M_{2}$ covering exactly $n-\delta-2$ vertices of $G_{1}$ that both avoid vertices of $Z\cup V_{\delta}\cup W$. We let $G_{2}$ be obtained from $G_{1}+M_{1}$ by adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup V(M_{1}))$. Then we have $d_{G_{2}}(x)=\delta+2$ and $d_{G_{2}}(v)=\delta+1$ for all $v\in V(G_{2})$ with $v\neq x$. Now we let $G_{3}$ be obtained from $G_{2}+M_{2}$ by adding a new vertex $y$ and all the edges joining $y$ and each vertex from $V(G_{2})\setminus(V(M_{2})\cup\{x\})$. The graph $G_{3}$ is $(\delta+2)$-regular and so has a Hamilton decomposition by Theorem 2.6. Deleting all the edges of $G_{3}$ incident with $x$ or $y$ and all the edges of $M_{1}\cup M_{2}$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+2)/2$ linear forests. Thus $\operatorname{la}(G)\leq g/2+(\delta+2)/2=(\Delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired.
Case 2: $g$ is odd.
Then we have $d_{G_{1}}(v)=\delta-1$ for each $v\in V_{\Delta}\setminus Z$, and $d_{G_{1}}(v)=\delta$ for each $v\in V_{\delta}\cup W\cup Z$. Thus $\Delta(G_{1})=\delta$.
If $\delta$ is even, as $G_{1}$ has precisely $n-(g+1)-|V_{\delta}|-|W|$ vertices of odd degrees, which are all $\delta-1$, we know that $n-(g+1)-|V_{\delta}|-|W|$ is even. As $\overline{G}_{1}$ has a matching covering at least $n-|V_{\delta}|-|Z|-1$ vertices of $G_{1}$ excluding vertices of $Z\cup W\cup V_{\delta}$, it has a matching $M$ covering exactly $n-(g+1)-\delta-|V_{\delta}|-|W|$ vertices not including any vertex of $Z\cup W\cup V_{\delta}$. We let $G_{2}$ be obtained from $G_{1}+M$ be adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup W\cup V_{\delta}\cup V(M))$. The graph $G_{2}$ is $\delta$-regular and so has a Hamilton decomposition by Theorem 2.6. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $\delta/2$ linear forests. Thus $\operatorname{la}(G)\leq(g+1)/2+\delta/2=(\Delta+1)/2$, as desired.
Next, consider that $\delta$ is odd. Then $G_{1}$ has precisely $g+1+|V_{\delta}|+|W|$ vertices of odd degrees, which are all $\delta$, it follows that $g+1+|V_{\delta}|+|W|$ is even. Thus if $n$ is even, then $n-(g+1)-|V_{\delta}|-|W|$ is even. We proceed exactly the same way above in getting a $\delta$-regular graph $G_{2}$. By Theorem 1.2, we have $\operatorname{la}(G_{2})\leq(\delta+1)/2$. Thus $\operatorname{la}(G)\leq(g+1)/2+(\delta+1)/2=\lceil(\Delta+1)/2\rceil$, as desired.
Thus we assume that $n$ is odd. As all the edges in $u_{2\lceil g/2\rceil}P_{1}u_{p-1}$ are contained in $\overline{G}_{1}$, and $V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})=V(G)\setminus(Z\cup V_{\delta}\cup W)$, we let $M_{1}$ be a matching of $\overline{G}_{1}$ covering all the vertices of $V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})\setminus\{u_{p-1}\}$. Let $G_{2}$ be obtained from $G_{1}$ by adding all the edges of $M_{1}$ and the edge $u_{p-1}u_{p}$ from $\overline{G}_{1}$. Then in $G_{2}$, all vertices other than $u_{p}$ have degree $\delta$ and $u_{p}$ has degree $\delta+1$. The graph $\overline{G}_{2}$ has one single vertex of degree $n-1-(\delta+1)$ and all other vertices of degree $n-1-\delta$. By Lemma 2.8, $\overline{G}_{2}$ has a matching covering at least $n-n/(n-\delta)-3$ vertices avoiding the vertex $u_{p}$. As $n$ is odd and $\delta$ is odd, we have $\delta\leq n-2$. As a consequence, $n-n/(n-\delta)-3\geq n/2-3>n-1-(\delta+1)$. Thus $\overline{G}_{2}$ has a matching $M_{2}$ covering exactly $n-1-(\delta+1)$ vertices excluding the vertex $u_{p}$. We let $G_{3}$ be obtained from $G_{2}+M_{2}$ by adding a new vertex $x$ and all the edges joining $x$ and vertices from $V(G_{2})\setminus(V(M_{2})\cup\{u_{p}\})$. The graph $G_{3}$ is $(\delta+1)$-regular and so has a Hamilton decomposition by Theorem 2.6. Deleting all the edges of $G_{3}$ incident with $x$ and all the edges of $M_{1}\cup M_{2}$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+1)/2$ linear forests. Thus $\operatorname{la}(G)\leq(g+1)/2+(\delta+1)/2=(\Delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired. ∎
By Claims 3.1 and 3.3, we assume that $g\leq 2$ or $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$. The following two cases together cover these two situations.
Case 1: $|U|=0$ ($|U|=0$ includes the cases that $g\leq 2$ or $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$ but $g<\eta n$).
The assumption of this case implies that $g<\eta n$. Since $d_{G}(v_{1})\leq\ldots\leq d_{G}(v_{n})$, we have $\operatorname{def}_{G}(v_{1},\Delta)\geq\ldots\geq\operatorname{def}_{G}(v_{n},\Delta)$. If $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is even and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, applying Claim 3.2 with $d=\Delta$, we can reduce $G$ into a regular graph $G^{*}$ with degree $\Delta-2k$ by taking out $k\leq\varepsilon n/8$ edge-disjoint linear forests. By Theorem 1.2, we have $\operatorname{la}(G^{*})\leq\lceil(\Delta-2k+1)/2\rceil$. Thus $\operatorname{la}(G)\leq k+\lceil(\Delta-2k+1)/2\rceil=\lceil(\Delta+1)/2\rceil$.
Thus we assume that $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd or $\operatorname{def}_{G}(v_{1},\Delta)>\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. If $\operatorname{def}_{G}(v_{1},\Delta)>\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, then by direct calculations, we know that $v_{1}$ is the only vertex of minimum degree in $G$ and $|V_{\delta+1}|\leq 1$. If $g\geq 3$, then we have $|W|\leq 1$ by Claim 3.1. Thus we assume that $g\leq 2$. As $\operatorname{def}_{G}(v_{1},\Delta)>\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, it follows that $\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)\leq 1$. This again implies that $|W|\leq 1$. Now Claim 3.3 applied to $G$ gives $\operatorname{la}(G)\leq(\Delta+1)/2$.
Thus we suppose that $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. Since $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)=n\Delta-2e(G)$, it follows that both $n$ and $\Delta$ are odd. Since $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, we know $\operatorname{def}_{G}(v_{3},\Delta)>0$. Let $P$ be a Hamilton $(v_{1},v_{n})$-path in $G$ by Theorem 2.4. Let $G_{1}=G-E(P)$. Then we have $d_{G_{1}}(w)=d_{G}(w)-2$ for $w\in V(G_{1})$ with $w\not\in\{v_{1},v_{n}\}$, $d_{G_{1}}(v_{1})=\delta-1$ and $d_{G_{1}}(v_{n})=\Delta-1$. Thus $\Delta(G_{1})=\Delta-1$ and $\delta-2\leq\delta(G_{1})\leq\delta-1$.
For every vertex $w\in V_{\Delta}$, by the definition, we have $\operatorname{def}_{G_{1}}(w,\Delta-2)=\operatorname{def}_{G}(w,\Delta)=0$; for every vertex $w\in V(G)\setminus V_{\Delta}$ with $w\neq v_{1}$, we have $\operatorname{def}_{G_{1}}(w,\Delta-2)=\operatorname{def}_{G}(w,\Delta)$, and $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)=\operatorname{def}_{G}(v_{1},\Delta)-1$. Therefore, we get $\sum_{i=1}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)=\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)-1$ and so $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta-2)$ is even. If we have $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)\geq\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$, then we have $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)<\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)=\sum_{i=2}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)$. Thus we suppose $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)<\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$. Then we have $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)=\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+1$ as $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+1=\operatorname{def}_{G}(v_{1},\Delta)\geq\operatorname{def}_{G}(v_{2},\Delta)=\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$. In this case, $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$ is the largest value among $\operatorname{def}_{G_{1}}(v_{i},\Delta-2)$ for all $i\in[1,n]$. Since $\operatorname{def}_{G_{1}}(v_{3},\Delta-2)=\operatorname{def}_{G}(v_{3},\Delta)>0$, we have $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)\leq\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+\sum_{i=3}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)$.
Applying Claim 3.2 with $G_{1}$ in the place of $G$ ($\delta(G_{1})\geq\delta(G)-2$, and the value $\delta(G)-2$ works essentially the same as $\delta(G)$ when $n$ is sufficiently large) and with $d=\Delta-2$, we can reduce $G_{1}$ into a graph $G^{*}$ by taking out $k\leq\varepsilon n/8$ edge-disjoint linear forests such that for any $v\in V(G_{1})$, we have $d_{G^{*}}(v)=d_{G_{1}}(v)-2k+\operatorname{def}_{G_{1}}(v,\Delta-2)$. For any $i\in[1,n-1]$, as $d_{G_{1}}(v_{i})\leq\Delta-2$, it follows that $d_{G_{1}}(v_{i})+\operatorname{def}_{G_{1}}(v_{i},\Delta-2)=\Delta-2$. Thus all vertices except $v_{n}$ of $G^{*}$ have degree $\Delta-2-2k$ and $v_{n}$ has degree $\Delta-1-2k$.
Applying Lemma 2.8, $\overline{G^{*}}$ has a matching excluding the vertex $v_{n}$ and covering at least $n-\frac{n}{n-(\Delta-2-2k)}-3$ vertices of $G^{*}$. Since $n$ is odd and $\Delta$ is odd, we have $\Delta\leq n-2$. Thus $n-\frac{n}{n-(\Delta-2-2k)}-3>n-1-((\Delta-2-2k)+1)$. Thus $\overline{G^{*}}$ has a matching $M$ covering exactly $n-1-((\Delta-2-2k)+1)$ vertices excluding the vertex $v_{n}$. We let $G_{2}$ be obtained from $G^{*}+M$ by adding a new vertex $x$ and all the edges joining $x$ and vertices from $V(G^{*})\setminus(V(M)\cup\{v_{n}\})$. The graph $G_{2}$ is $((\Delta-2-2k)+1)$-regular and so has a Hamilton decomposition by Theorem 2.6. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G^{*}$ into $((\Delta-2-2k)+1)/2$ linear forests. Thus $\operatorname{la}(G)\leq 1+k+((\Delta-2-2k)+1)/2=(\Delta+1)/2$, as desired.
Case 2: $|U|\geq\eta n$ ($|U|\geq\eta n$ includes the case that $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$ and $g\geq\eta n$).
If $\Delta$ is odd, we delete a perfect matching $M$ from $G$ if $n$ is even, and we delete a matching covering all vertices of $V(G)\setminus\{v_{1}\}$ if $n$ is odd, where recall that we assumed $d_{G}(v_{1})=\delta$ and $M$ exists as $G$ is Hamilton-connected by Theorem 2.4. Thus in the following, we let $\Delta^{*}=\Delta$ and $G^{*}=G$ if $\Delta$ is even, and $\Delta^{*}=\Delta-1$ and $G^{*}=G-M$ if $\Delta$ is odd. Note that $\sum_{i=1}^{n}\operatorname{def}_{G^{*}}(v_{i},\Delta^{*})\geq(\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta))-1$. We will add vertices to $G^{*}$ and construct a simple $\Delta^{*}$-regular graph $H$ in three steps such that $G^{*}\subseteq H$, and $H$ is a robust $(\eta^{2},2\eta)$-expander.
Let $\operatorname{def}(G^{*})=\sum_{i=1}^{n}\operatorname{def}_{G^{*}}(v_{i},\Delta^{*})$. We add a set $W$ of new vertices to $G^{*}$ such that $|W|$ is the smallest integer satisfying the following properties:
  1. (1)
    $|W|\equiv n\ (\mbox{mod}\ 2)$;
  2. (2)
    $|W|\geq\Delta^{*}-\lfloor\frac{{\rm def}(G^{*})}{|W|}\rfloor+1$.
Note that when $|W|=\Delta^{*}+1$ or $|W|=\Delta^{*}+2$, the conditions above are satisfied. Thus a smallest integer $|W|$ satisfying the conditions above exists and $|W|\leq\Delta^{*}+2\leq n+1$. Since $|W|\geq\Delta^{*}-\frac{{\rm def}(G^{*})}{|W|}+1$, solving the corresponding quadratic inequality involving $|W|$ gives $|W|\geq\frac{(\Delta^{*}+1)+\sqrt{(\Delta^{*}+1)^{2}-4{\rm def}(G^{*})}}{2}\geq\frac{\Delta^{*}+1}{2}$. Note that $(\Delta^{*}+1)^{2}-4\operatorname{def}(G^{*})\geq 0$: we have $\operatorname{def}(G^{*})<(\Delta^{*}-\delta(G^{*}))n/2$ by $|V_{\Delta}|>n/2$ from (2) (the set of maximum degree vertices of $G$ and $G^{*}$ keep the same) and so
$(\Delta^{*}+1)^{2}-4\operatorname{def}(G^{*})$$\geq$ $(\Delta^{*}+1)^{2}-2\Delta^{*}n+2\delta(G^{*})n$
$>$ $(\Delta^{*}+1)^{2}-2\Delta^{*}n+n^{2}>(\Delta^{*}-n)^{2}\geq 0.$
Suppose that ${\rm def}(G^{*})\equiv\ell\ (\mbox{mod}\ |W|)$, where $\ell\in[0,|W|-1]$. Let $d_{1}=\ldots=d_{\ell}=\Delta^{*}-\lceil\frac{{\rm def}(G^{*})}{|W|}\rceil=d-1$, and $d_{\ell+1}=\ldots=d_{|W|}=\Delta^{*}-(\lceil\frac{{\rm def}(G^{*})}{|W|}\rceil-1)=d$. Since $|W|\equiv n\ (\mbox{mod}\ 2)$, we know that
$\sum\limits_{i=1}^{|W|}d_{i}=\Delta^{*}|W|-{\rm def}(G^{*})\equiv\Delta^{*}n-{\rm def}(G^{*})=2e(G^{*})\equiv 0\ (\mbox{mod}\ 2).$
Furthermore, $d=|W|\geq 2$. Let $W=\{w_{1},\ldots,w_{|W|}\}$. By Lemma 2.2, there is a graph $R$ on $W$ such that $d_{R}(w_{i})=d_{i}$ for each $i\in[1,|W|]$. We now add ${\rm{def}}(G^{*})$ edges joining vertices from $V(G^{*})\setminus V_{\Delta^{*}}$ to $W$. We list these ${\rm{def}}(G^{*})$ edges as $e_{1},e_{2},\ldots,e_{{\rm{def}}(G^{*})}$ and add them one by one such that
  1. (1)
    For each vertex $v\in V(G^{*})\setminus V_{\Delta^{*}}$, the ${\rm def}_{G^{*}}(v)$ edges joining $v$ with the vertices of $W$ are listed consecutively in the ordering above.
  2. (2)
    For each $w_{i}\in W$, $w_{i}$ is incident with $\Delta^{*}-d_{i}$ edges for each $i\in[1,|W|]$.
The resulting multigraph is denoted by $H$. Then $H$ is $\Delta^{*}$-regular by the above construction. The multigraph $H$ is actually simple, since $|W|\geq(\Delta^{*}+1)/2>\Delta^{*}-\delta(G^{*})$ by $\delta(G^{*})\geq(1+\varepsilon)n/2-1$. Let $|V(H)|=n+|W|=n^{\prime}$. As $|W|\leq n+1$ and $\Delta^{*}\geq\delta+\eta n-1>n/2$, it follows that $\Delta^{*}>n^{\prime}/4$. In the rest part of the proof, we mainly show that $H$ is a robust $(\eta^{2},2\eta)$-expander. Let $S\subseteq V(H)$ with $2\eta n^{\prime}\leq|S|\leq(1-2\eta)n^{\prime}$. We show that $|RN_{H}(S)|\geq|S|+\eta^{2}n^{\prime}$.
Consider first that $|S\cap V(G^{*})|\geq\eta n/2$. Since $G^{*}$ is a robust $(5\eta^{2},\frac{\eta}{2})$-expander by Lemma 2.10, we have $|RN_{G^{*}}(S)|\geq|S\cap V(G^{*})|+5\eta^{2}n$. If $|S\cap W|\leq\eta^{2}n^{\prime}+3$, then
$|RN_{H}(S)|$ $\geq|RN_{G^{*}}(S)|\geq|S\cap V(G^{*})|+|S\cap W|+2.5\eta^{2}n$
$\geq|S|+\eta^{2}n^{\prime},$
as $|W|\leq n+1$ and so $n^{\prime}\leq 2n+1$. Thus we assume that $|S\cap W|>\eta^{2}n^{\prime}+3$. Then as $\delta(R)=d-1>|W|-4$ by the minimality of $|W|$ and the definition of $d$, we get $W\subseteq RN_{H}(S)$. Therefore,
$|RN_{H}(S)|\geq|RN_{G^{*}}(S)|+|W\cap S|\geq|S\cap V(G^{*})|+5\eta^{2}n+|W\cap S|\geq|S|+\eta^{2}n^{\prime}.$
Consider then that $|S\cap V(G^{*})|<\eta n/2$. As $|S|\geq 2\eta n$, we have that $|S\cap W|>3\eta n/2$. Again, as $\delta(R)=d-1>|W|-4$ by the minimality of $|W|$ and the definition of $d$, we have $W\subseteq RN_{H}(S)$. If $|W\backslash S|\geq\eta^{2}n^{\prime}+\eta n/2$, then
$|RN_{H}(S)|\geq|W|=|S|-|S\cap V(G^{*})|+|W\setminus S|\geq|S|-\eta n/2+\eta^{2}n^{\prime}+\eta n/2=|S|+\eta^{2}n^{\prime}.$
Therefore, we have $|W\backslash S|<\eta^{2}n^{\prime}+\eta n/2$. Since every vertex from $U$ has in $H$ at least $\eta n$ neighbors from $W$, we then know that $U\subseteq RN_{H}(S)$ and so
$|RN_{H}(S)|\geq|W|+|U|\geq|W\cap S|+|V(G^{*})\cap S|+\eta n/3\geq|S|+\eta^{2}n^{\prime}.$
Thus $H$ is a $\Delta^{*}$-regular robust $(\eta^{2},2\eta)$-expander, where $\Delta^{*}\geq n^{\prime}/4$ is even. As $\eta\ll\tau^{*}$ by (1), $H$ is a robust $(\eta^{2},2\eta)$-expander implies that $H$ is a robust $(\eta^{2},\tau^{*})$-expander. As $n$ is taken to be at least $N^{*}_{0}$ as defined in Theorem 2.9, the theorem implies that $H$ has a Hamilton decomposition into $\Delta^{*}/2$ Hamilton cycles. By deleting the edges incident with vertices of $W$ from the decomposition, we obtain a decomposition of $G^{*}$ into $\Delta^{*}/2$ linear forests. If $\Delta=\Delta^{*}$, we get $\operatorname{la}(G)\leq\Delta/2$. Otherwise, we get $\operatorname{la}(G)\leq 1+(\Delta-1)/2=(\Delta+1)/2$. The proof of Theorem 1.3 is now completed. ∎