### v2

# Linear arboricity of robust expanders

Abstract
In 1980, Akiyama, Exoo, and Harary conjectured that any graph $G$ can be decomposed into at most $\lceil(\Delta(G)+1)/2\rceil$ linear forests. We confirm the conjecture for robust expanders of linear minimum degree. As a consequence, the conjecture holds for dense quasirandom graphs of linear minimum degree as well as for large $n$-vertex graphs with minimum degree arbitrarily close to $n/2$ from above.
Keywords: Linear forest; linear arboricity; robust expander; Hamilton decomposition
1 Introduction
Unless otherwise stated explicitly, graphs in this paper are simple and finite. A linear forest is a graph consisting of vertex-disjoint paths. The linear arboricity of a graph $G$, denoted $\operatorname{la}(G)$, is the smallest number of linear forests its edges can be partitioned into. This notion was introduced by Harary [15] in 1970 as one of the covering invariants of graphs, and has been studied quite extensively since then. The linear arboricity of any graph $G$ of maximum degree $\Delta$ is known to be at least $\lceil\Delta/2\rceil$ as the edges incident with a maximum degree vertex need to be covered by at least $\lceil\Delta/2\rceil$ distinct linear forests. On the other hand, $\operatorname{la}(G)$ is conjectured to be at most $\lceil(\Delta+1)/2\rceil$. The conjecture, known as the Linear Arboricity Conjecture, was proposed by Akiyama, Exoo and Harary [1] in 1980.
Conjecture 1.1 (Linear Arboricity Conjecture).
Every graph $G$ satisfies $\operatorname{la}(G)\leq\lceil(\Delta(G)+1)/2\rceil$.
The conjecture would determine the linear arboricity exactly for graphs of odd maximum degree, since in that case the upper and lower bounds coincide. For graphs of even maximum degree it would imply that the linear arboricity must be one of only two possible values. However, determining the exact value among these two choices is NP-complete [20].
In terms of general bounds on $\operatorname{la}(G)$ for graphs $G$ with maximum degree $\Delta$, the first result of this type was obtained by Alon in 1988 [2], which states that $\operatorname{la}(G)\leq\Delta/2+O\left(\frac{\Delta\log\log\Delta}{\log\Delta}\right)$. The error term was improved to $O(\Delta^{2/3}\log^{1/3}\Delta)$ by Alon and Spencer in 1992 [3], and has been further improved to $O(\Delta^{0.661})$ by Ferber, Fox and Jain [10], and then to $O(\Delta^{1/2}\log^{4}\Delta)$ by Lang and Postle [18]. Very recently, Christoph, Draganić, Girão, Hurley, Michel, and Müyesser [6] made a significant improvement by reducing the error term to $O(\log n)$, which is an exponential improvement over the previous best error term if $\Delta=\Omega(n^{\varepsilon})$ for any given $0<\varepsilon\leq 1$. The Linear Arboricity Conjecture was well studied for classes of graphs including degenerate graphs, graphs with small maximum degree, and planar graphs. For example, see [5, 7, 9, 13, 19, 24, 26, 27]. The conjecture was confirmed for random graphs $G_{n,p}$ with constant probability $p$ by Glock, Kühn, and Osthus [12] in 2016, and for the range $\frac{C\log n}{n}\leq p\leq n^{-2/3}$ by Draganić, Glock, Munhá Correia, and Sudakov in 2025, where $C$ is a constant [8]. Glock, Kühn, and Osthus [12] also verified the conjecture for dense quasirandom graphs in which the difference between the maximum and minimum degrees is small. Here, we use the following notion of quasirandomness, which is a one-sided version of $\varepsilon$-regularity. Let $0<\varepsilon,p<1$. A graph $G$ on $n$ vertices is called lower-$(p,\varepsilon)$-regular if $e_{G}(S,T)\geq(p-\varepsilon)|S||T|$ for all disjoint $S,T\subseteq V(G)$ with $|S|,|T|\geq\varepsilon n$.
The Linear Arboricity Conjecture holds for regular robust expanders of linear minimum degree, which we define now. Given $0<\nu\leq\tau<1$, we say that a graph $G$ on $n$ vertices is a robust $(\nu,\tau)$-expander, if for all $S\subseteq V(G)$ with $\tau n\leq|S|\leq(1-\tau)n$ the number of vertices that have at least $\nu n$ neighbors in $S$ is at least $|S|+\nu n$. The $\nu$-robust neighborhood $RN_{G}(S)$ is the set of all those vertices of $G$ which have at least $\nu n$ neighbors in $S$. Note that for $0<\nu^{\prime}\leq\nu\leq\tau\leq\tau^{\prime}<1$, any robust $(\nu,\tau)$-expander is also a robust $(\nu^{\prime},\tau^{\prime})$-expander. The theorem below by Kühn and Osthus [17] is profound and also is a powerful tool that has been used to establish other results.
Theorem 1.2 ([17]).
For every $\alpha>0$ there exists $\tau=\tau(\alpha)>0$ such that for every $\nu>0$ there exists an integer $N_{0}=N_{0}(\alpha,\nu,\tau)$ for which the following holds. Suppose that
  1. (i)
    $G$ is an $r$-regular graph on $n\geq N_{0}$ vertices, where $r\geq\alpha n$ is even;
  2. (ii)
    $G$ is a robust $(\nu,\tau)$-expander.
Then $G$ has a Hamilton decomposition. Moreover, this decomposition can be found in polynomial time in $n$.
As an application of Theorem 1.2, the Linear Arboricity Conjecture holds for regular robust expanders of linear minimum degree, as stated below. The proof follows exactly the same proof of Theorem 1.4 in [12], and we omit it here.
Theorem 1.3.
Let $n,r\in\mathbb{N}$ and suppose that $0<1/n\ll\nu\leq\tau,\alpha<1$. If $G$ is a robust $r$-regular $(\nu,\tau)$-expander on $n$ vertices with $r\geq\alpha n$, then $\operatorname{la}(G)\leq\lceil\frac{r+1}{2}\rceil$.
Using Theorem 1.2, Glock, Kühn, and Osthus [12] in 2016 also proved the Linear Arboricity Conjecture for regular graphs of large degree.
Theorem 1.4 ([12]).
There exists $n_{0}\in\mathbb{N}$ such that every $d$-regular graph $G$ on $n\geq n_{0}$ vertices with $d\geq\left\lfloor\frac{n-1}{2}\right\rfloor$ has a decomposition into $\left\lceil\frac{d+1}{2}\right\rceil$ linear forests.
In this paper, we investigate the Linear Arboricity Conjecture for robust expanders of linear minimum degree that are not necessarily regular and obtain the following result.
Theorem 1.5.
For every $\alpha>0$ there exists $\tau=\tau(\alpha)>0$ such that for every $\nu>0$ there exists an integer $N_{0}=N_{0}(\alpha,\nu,\tau)$ for which the following holds. If $G$ is a robust $(\nu,\tau)$-expander on $n\geq N_{0}$ vertices with $\delta(G)\geq\alpha n$, then $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$.
Note that Conjecture 1.1 is equivalent to its restriction on regular graphs, as any graph $G$ with maximum degree $\Delta$ can be embedded in a $\Delta$-regular graph $H$. However, Theorem 1.5 does not follow from Theorem 1.3. For example, for sufficiently large $n$, we let $G$ be obtained from $K_{n-3}$ and $K_{3}$ by adding edges joining each vertex of the $K_{3}$ and about (n3)/3(n-3)/3 distinct vertices of the $K_{n-3}$ such that the resulting graph has maximum degree n3n-3. It is clear that the graph is a robust expander. However, any (n3)(n-3)-regular supergraph $H$ of $G$ has at least $n+(n-3-\frac{n+3}{3})$ vertices and has a cut of size 3. The graph is not a robust $(\nu,\tau)$-expander for any $0<\nu\leq\tau\leq 2/5$.
It was shown independently by Glock, Kühn, and Osthus [12] and Kühn and Osthus [17] that dense quasirandom graphs and graphs with large minimum degree are robust expanders.
Lemma 1.6 ([12]).
Let $0<1/n_{0}\ll\varepsilon\ll\tau,p<1$ and suppose $G$ is a lower-$(p,\varepsilon)$-regular graph on $n\geq n_{0}$ vertices. Then $G$ is a robust $(\varepsilon,\tau)$-expander.
Lemma 1.7 ([17]).
Suppose that $0<\nu\leq\tau\leq\varepsilon<1$ are such that $\varepsilon\geq 4\nu/\tau$. Let $G$ be a graph on $n$ vertices with minimum degree $\delta(G)\geq(1+\varepsilon)n/2$. Then $G$ is a robust $(\nu,\tau)$-expander.
As a consequence of Theorem 1.5 and Lemmas 1.6 and 1.7, we have the following corollaries.
Corollary 1.8.
For any given $0<\varepsilon,p,\alpha<1$, there exists $n_{0}\in\mathbb{N}$ for which the following statement holds. If $G$ is a lower-$(p,\varepsilon)$-regular graph on $n\geq n_{0}$ vertices with $\delta(G)\geq\alpha n$, then $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$.
Corollary 1.9.
For any given $0<\varepsilon<1$, there exists $n_{0}\in\mathbb{N}$ for which the following statement holds. If $G$ is a graph on $n\geq n_{0}$ vertices with $\delta(G)\geq\frac{1}{2}(1+\varepsilon)n$, then $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$.
Note again that Corollary 1.9 does not follow from Theorem 1.4 as passing to a regular supergraph may significantly increase the number of vertices. Indeed, we now construct a graph $G$ of maximum degree $\Delta$ such that any $\Delta$-regular supergraph $H$ of $G$ satisfies $\Delta<|V(H)|/2$. Choose $n$ such that both $\frac{1}{2}n$ and $\varepsilon n$ are even integers. We let $G_{1}$ be a $3\varepsilon n$-regular graph on $\frac{1}{2}n-\varepsilon n$ vertices, $G_{2}$ be a $5\varepsilon n$-regular graph on another $\frac{1}{2}n-\varepsilon n$ vertices, and $G_{3}$ be another complete graph on $2\varepsilon n$ vertices. Then the graph $G$ is obtained from $G_{1},G_{2},G_{3}$ by adding all edges between $G_{1}$ and $G_{2}$, and all edges between $G_{1}$ and $G_{3}$. Now we have $\delta(G)=\frac{1}{2}n-\varepsilon n+2\varepsilon n-1=\frac{1}{2}n+\varepsilon n-1$, and $\Delta:=\Delta(G)=\frac{1}{2}n-\varepsilon n+3\varepsilon n+2\varepsilon n=\frac{1}{2}n+4\varepsilon n$. Also, all vertices of $V(G_{1})\cup V(G_{2})$ are of maximum degree in $G$. Thus we have $\sum_{v\in V(G)}(\Delta-d_{G}(v))=2\varepsilon n(3\varepsilon n+1)$. Since all vertices of minimum degree in $G$ form a clique, and $\Delta(G)-\delta(G)=3\varepsilon n+1$, it follows that any $\Delta$-regular supergraph $H$ of $G$ has at least $n+3\varepsilon n+1$ vertices. On average, one vertex $v$ from $V(H)\setminus V(G)$ is adjacent in $H$ to at most $2\varepsilon n$ vertices from $V(G)$. As $H$ is $\Delta$-regular, $v$ has at least $\Delta-2\varepsilon n$ neighbors in $H$ from $V(H)\setminus V(G)$. Thus $H$ has more than $n+\Delta-2\varepsilon n>\frac{3}{2}n$ vertices. As a consequence, we have $\Delta<\frac{1}{3}(1+8\varepsilon)|V(H)|$, where $\frac{1}{3}(1+8\varepsilon)|V(H)|<\frac{1}{2}|V(H)|$ if $\varepsilon<\frac{1}{16}$.
Our proof of Theorem 1.5, which relies on Theorem 1.2, will be given in Section 3 after some preliminaries in the next section.
2 Notation and Preliminaries
Let $G$ be a graph and $S\subseteq V(G)$. We let $n(G)=|V(G)|$ and $e(G)=|E(G)|$. The subgraphs of $G$ induced on $S$ and GSG-S are respectively denoted by $G[S]$ and GSG-S. For $v\in V(G)$, we write GvG-v for $G-\{v\}$. For two disjoint subsets $S,T$ of $V(G)$, we denote by $E_{G}(S,T)$ the set of edges in $G$ with one endvertex in $S$ and the other in $T$, and let $e_{G}(S,T)=|E_{G}(S,T)|$. If $F\subseteq E(G)$, then GFG-F is obtained from $G$ by deleting all edges in $F$. If $F$ is a submultiset of edges from the complement $\overline{G}$, then $G+F$ is obtained from $G$ by adding the edges of $F$ to $G$ ($G+F$ may become a multigraph). For two integers $p$ and $q$, let $p,q]=\{i\in\mathbb{Z}\mid p\leq i\leq q\}$.
We will make use of the following notation: $0<a\ll b\leq 1$. Precisely, if we claim that a statement is true whenever $0<a\ll b\leq 1$, then this means that there exists a non-decreasing function $f:(0,1]\rightarrow(0,1]$ such that the statement holds for all $0<a\ll b\leq 1$ satisfying $a\leq f(b)$. Intuitively, this means that for every $b>0$, the statement is true provided that $a$ is sufficiently small compared to $b$.
2.1 Degree sequences and number of maximum degree vertices
Let $n\geq 2$ be an integer. A sequence of nonnegative integers listed in non-increasing order $(d_{1},\ldots,d_{n})$ is graphic if there is a graph $G$ such that the degree sequence of $G$ is $(d_{1},\ldots,d_{n})$. In this case, we say that $G$ realizes $(d_{1},\ldots,d_{n})$.
The first result below by Hakimi [14] gives a sufficient and necessary condition for a sequence of integers to be the degree sequence of a multigraph. The second result by the second author [22] will be used to construct a supergraph of a given graph $G$ by adding additional vertices.
Theorem 2.1 ([14]).
Let $0\leq d_{n}\leq\ldots\leq d_{1}$ be integers. Then there exists a multigraph $G$ on vertices $v_{1},v_{2},\ldots,v_{n}$ such that $d_{G}(v_{i})=d_{i}$ for all $i\in[1,n]$ if and only if $\sum\limits_{i=1}^{n}d_{i}$ is even and $\sum\limits_{i=2}^{n}d_{i}\geq d_{1}$.
Lemma 2.2 ([22]).
Let $n$ and $d$ be positive integers such that $n\geq d+1\geq 3$. Suppose $(d_{1},\ldots,d_{n})$ is a sequence of positive integers with $d_{1}=\ldots=d_{t}=d$ and $d_{t+1}=\ldots=d_{n}=d-1$ and $\sum\limits_{i=1}^{n}d_{i}$ even, where $t\in[1,n]$. Then $(d_{1},\ldots,d_{n})$ is graphic. Furthermore, the sequence can be realized in polynomial time in $n$.
We will also need the following result by Plantholt and the second author [21] which gives a lower bound on the number of maximum degree vertices if we assume that all vertices of degree less than maximum degree form a clique.
Lemma 2.3 ([21]).
Let $G$ be a graph on $n$ vertices such that all vertices of degree less than $\Delta(G)$ are mutually adjacent in $G$. Then $G$ has more than $n/2$ vertices of maximum degree.
2.2 Matchings in almost-regular graphs
Let $G$ be a graph. A matching $M$ in $G$ is a set of vertex-disjoint edges. The set of vertices covered or saturated by $M$ is denoted by $V(M)$. We say that $M$ is perfect if $V(M)=V(G)$. The graph $G$ is factor-critical if $G$ does not have a perfect matching but GvG-v has one for any $v\in V(G)$.
For a subset $X\subseteq V(G)$, we denote by o(GX)o(G-X) the number of components of GXG-X of odd order. Define ${\rm df}(G)=\max_{X\subseteq V(G)}(o(G-X)-|X|)$. By applying Tutte’s Theorem [23] to the graph obtained from $G$ by adding ${\rm df}(G)$ vertices and joining all edges between $V(G)$ and these new vertices, Berge [4] observed that the maximum size of a matching in $G$ is $\frac{1}{2}(n(G)-{\rm df}(G))$. West [25] gave a short proof of this result. To state his result, we need some more definitions.
For $X\subseteq V(G)$, denote by $Y$ the set of components of GXG-X. We define an auxiliary bipartite multigraph $B(X)$ by contracting each component of GXG-X to a single vertex and deleting edges within $X$. Thus for any vertex $x\in X$ and any $y\in V(B(X))\setminus X$ such that $y$ is contracted from a component $D_{y}$ of $Y$, there are exactly $e_{G}(\{x\},V(D_{y}))$ edges between $x$ and $y$ in $B(X)$.
Lemma 2.4.
Let $G$ be a multigraph and $X\subseteq V(G)$ be maximal with $o(G-X)-|X|={\rm df}(G)$. Then the following statements hold.
  1. (1)
    ([25, Lemma 2]) Every component of GXG-X has odd order and is factor-critical.
  2. (2)
    ([25, Lemma 3]) The bipartite multigraph $B(X)$ has a matching covering $X$.
  3. (3)
    The bipartite multigraph $B(X)$ has a matching covering $X$ and all vertices of degree at least $d$ from $V(B(X))\setminus X$ in $B(X)$, where $d=\max\{d_{B(X)}(x)\mid x\in X\}$.
Proof.
Statements (1)-(2) were proved by West in [25].
We prove statement (3). By (2), $B(X)$ has a matching $M_{1}$ covering $X$. Let $Y=V(B(X))\setminus X$, $d=\max\{d_{B(X)}(x)\mid x\in X\}$, $Y_{d}=\{y\in Y\mid d_{B(X)}(y)\geq d\}$. By applying Hall’s Theorem to $B(X)[X,Y_{d}]$, $B(X)[X,Y_{d}]$ and so $B(X)$ has a matching $M_{2}$ saturating all vertices from $Y_{d}$, where $B(X)[X,Y_{d}]$ is the induced subgraph of $B(X)$ with bipartitions as $X$ and $Y_{d}$. In each component $D$ of the subgraph of $B(X)$ induced on the symmetric difference of $M_{1}$ and $M_{2}$, there is a matching that saturates vertices in $V(D)\cap(X\cup Y_{d})$. The union of these matchings together with edges of $M_{1}\cap M_{2}$ gives a desired matching that covers all vertices of $X\cup Y_{d}$. ∎
A graph $G$ is almost $r$-regular for some integer $r\geq 0$ if exactly one vertex has degree $r+1$ and all others have degree $r$.
Lemma 2.5.
Let $G$ be an $n$-vertex almost $r$-regular graph for some integer $r\geq 0$. Then $\overline{G}$ has a matching covering at least $n-\frac{n}{n-r}-3$ vertices of $\overline{G}$ excluding the vertex of degree $r+1$ in $G$.
Proof.
Let $x\in V(G)$ be the vertex with $d_{G}(x)=r+1$. Note that $\overline{G}$ has n1n-1 vertices of degree (n1r)(n-1-r) and one vertex $x$ of degree n2rn-2-r. If $\overline{G}$ has a perfect matching, then we are done. Thus assume that $\overline{G}$ does not have a perfect matching. Let $X\subseteq V(\overline{G})$ be maximal such that $o(\overline{G}-X)-|X|={\rm df}(\overline{G})$. Then $\overline{G}-X$ has no even component by Lemma 2.4(1). For any odd component $D$ of $\overline{G}-X$ such that $x\not\in V(D)$ and $|V(D)|\leq n-r-1$, we have
$e_{\overline{G}}(V(D),X)$ $\geq$ $|V(D)|(n-r-1-(|V(D)|-1))$
$\geq$ $n-r-1,$
where the last inequality is obtained as the quadratic function |V(D)|(nr1(|V(D)|1))|V(D)|(n-r-1-(|V(D)|-1)) in $|V(D)|$ achieves its minimum at $|V(D)|=1$ or |V(D)|=nr1|V(D)|=n-r-1. There are at most $\frac{n}{n-r}$ odd components $D$ of $\overline{G}-X$ with $|V(D)|\geq n-r$, and so there are at most $\frac{n}{n-r}+1$ odd components $D$ of $\overline{G}-X$ such that $e_{\overline{G}}(V(D),X)\leq n-r-2$. As every vertex of $X$ has degree at most n1rn-1-r in $\overline{G}$, by Lemma 2.4(3), we know that $o(\overline{G}-X)-|X|\leq\frac{n}{n-r}+1$ and so ${\rm df}(\overline{G})\leq\frac{n}{n-r}+1$. Thus by Berge’s formula from [4] that the maximum size of a matching in a graph $H$ is $\frac{1}{2}(n(H)-{\rm df}(H))$, we conclude that $\overline{G}$ has a matching that covers at least $n-\frac{n}{n-r}-1$ vertices of $\overline{G}$, and so it has a matching that covers at least $n-\frac{n}{n-r}-3$ vertices of $\overline{G}$ excluding $x$. ∎
2.3 Robust expanders
The following lemma states that robust expansion is preserved when few edges are removed and/or few vertices are removed and/or added. The result is stated for robust outexpanders in [11], we list here its undirected counterpart.
Lemma 2.6 ([11, Lemma 4.2]).
Let $0<\varepsilon\leq\nu\leq\tau\leq 1$. Let $G$ be a robust $(\nu,\tau)$-expander on $n$ vertices.
  1. (a)
    If $G^{\prime}$ is obtained from $G$ by removing at most $\varepsilon n$ edges at each vertex, then $G^{\prime}$ is a robust $(\nu-\varepsilon,\tau)$-expander.
  2. (b)
    Suppose that $\tau\geq(1+2\tau)\varepsilon$. If $G^{\prime}$ is obtained from $G$ by adding or removing at most $\varepsilon n$ vertices, then $G^{\prime}$ is a robust $(\nu-\varepsilon,2\tau)$-expander.
Following [16], for a given integer $k\geq 2$, we say a graph $G$ is Hamilton $k$-linked if whenever $x_{1},y_{1},\ldots,x_{k},y_{k}$ are distinct vertices, there exist vertex-disjoint paths $P_{1},\ldots,P_{k}$ such that $P_{i}$ connects $x_{i},y_{i}$ for each $i\in[1,k]$ and such that together the paths $P_{1},\ldots,P_{k}$ cover all vertices of $G$. We will use the following result by Kühn, Lo, Osthus, and Staden from [16].
Lemma 2.7 ([16, Corollary 6.9(ii)]).
Let $n,k\in\mathbb{N}$ and suppose that $0<1/n\ll\nu\ll\tau\ll\alpha<1$ and $k\leq\nu^{4}n$. If $H$ is a robust $(\nu,\tau)$-expander on $n$ vertices with $\delta(H)\geq\alpha n$, then $H$ is Hamilton $k$-linked.
We next introduce the notion of a robust outexpander and cite a theorem of Girão, Granet, Kühn, Lo, and Osthus [11] on “almost-regular” graphs. The result will help us with linear forest decomposition when the graph is close to regular.
Let $D$ be a digraph on $n$ vertices. We write $xy$ for an edge which is directed from $x$ to $y$, and call $y$ an outneighbor of $x$ and $x$ an inneighbor of $y$. For $x\in V(D)$, the outdegree of $x$, denoted $d^{+}_{D}(x)$, is the number of edges $xy$ for all $y\in V(D)$; and the indegree of $x$, denoted $d^{-}_{D}(x)$, is the number of edges $yx$ for all $y\in V(D)$. Define the minimum semidegree of $D$ as $\delta^{0}(D)=\min\{d_{D}^{+}(x),d_{D}^{-}(x)\mid x\in V(D)\}$.
Let $0<\nu\leq\tau<1$ and $S\subseteq V(D)$. The $\nu$-robust outneighborhood $RN^{+}_{\nu,D}(S)$ of $S$ is the set of all those vertices $x$ of $D$ which have at least $\nu n$ inneighbors from $S$. The digraph $D$ is called a robust $(\nu,\tau)$-outexpander if
$|RN^{+}_{\nu,D}(S)|\geq|S|+\nu n$
for all $S\subseteq V(D)$ with $\tau n\leq|S|\leq(1-\tau)n$.
Let $\varepsilon,\delta>0$. We say $D$ is $(\delta,\varepsilon)$-almost regular if, for each $v\in V(D)$, $d^{+}_{D}(v)=(\delta\pm\varepsilon)n$ and $d^{-}_{D}(v)=(\delta\pm\varepsilon)n$. Similarly, if $D$ is a graph, then it is $(\delta,\varepsilon)$-almost regular if, for each $v\in V(D)$, it holds that $d_{D}(v)=(\delta\pm\varepsilon)n$.
Let $0<\varepsilon,p<1$. We say $D$ is an $(\varepsilon,p)$-robust $(\nu,\tau)$-outexpander if $D$ is a robust $(\nu,\tau)$-outexpander and, for any integer $k\geq\varepsilon n$, if $S\subseteq V(D)$ is a random subset of size $k$, then $D[S]$ is a robust $(\nu,\tau)$-outexpander with probability at least 1p1-p. By the definitions, we have the following fact.
Fact 1.
Let $D$ be an $(\varepsilon,p)$-robust $(\nu,\tau)$-outexpander. Then, for any $\varepsilon^{\prime}\geq\varepsilon$, $p^{\prime}\geq p$, $\nu^{\prime}\leq\nu$, and $\tau^{\prime}\geq\tau$, $D$ is an $(\varepsilon^{\prime},p^{\prime})$-robust $(\nu^{\prime},\tau^{\prime})$-outexpander.
The lemma below says that an induced subgraph on a subset with relatively large size of a robust outexpander is also a robust outexpander.
Lemma 2.8 ([11, Lemma 14.3]).
Let $0<\frac{1}{n}\ll\varepsilon\ll\nu^{\prime}\ll\nu,\tau,\alpha\ll 1$. Suppose $D$ is a robust $(\nu,\tau)$-outexpander on $n$ vertices satisfying $\delta^{0}(D)\geq\alpha n$. Then $D$ is an $(\varepsilon,n^{-2})$-robust $(\nu^{\prime},4\tau)$-outexpander.
We also need the following result by Kühn and Osthus.
Lemma 2.9 ([17, Lemma 3.2]).
Suppose that $0<1/n\ll\eta\ll\nu,\tau,\alpha,\lambda,1-\lambda<1$. Let $G$ be a digraph on $n$ vertices with $\delta^{0}(G)\geq\alpha n$ which is a robust $(\nu,\tau)$-outexpander. Then $G$ can be split into two edge-disjoint spanning subdigraphs $G_{1}$ and $G_{2}$ such that the following two properties hold.
  1. (i)
    $d^{+}_{G_{1}}(x)=(1\pm\eta)\lambda d^{+}_{G}(x)$ and $d^{-}_{G_{1}}(x)=(1\pm\eta)\lambda d^{-}_{G}(x)$ for every $x\in V(G)$.
  2. (ii)
    $G_{1}$ is a robust $(\lambda\nu/2,\tau)$-outexpander and $G_{2}$ is a robust $((1-\lambda)\nu/2,\tau)$-outexpander.
Let $D$ be a multidigraph on vertex set $V$. Let $L$ be a multiset consisting of paths on $V$ and isolated vertices. We use $V(L)$ to denote the set of vertices that are either isolated in $L$ or on the non-trivial paths contained in $L$, use $E(L)$ for the multiset of edges that are contained in paths of $L$, and for $v\in V(L)$, we use $d_{L}(v)$ to denote the number of edges from $E(L)$ that have $v$ as one endvertex. Let $F$ be a multiset of edges. We say $(L,F)$ is a layout if $F\subseteq E(L)$ and $E(L)\setminus F\neq\emptyset$. We define a layout for a multigraph exactly the same way as above. When $F$ is a multidigraph or multigraph, we also simply say that $(L,F)$ is a layout instead of $(L,E(F))$ being a layout.
Given an edge $uv$, we say that a $(u,v)$-path has shape $uv$. Let $(L,F)$ be a layout on $V$. A multidigraph $\mathcal{H}$ on $V$ is a spanning configuration of shape $(L,F)$ if $\mathcal{H}$ can be decomposed into internally vertex-disjoint paths $\{P_{e}\mid e\in E(L)\}$ such that each $P_{e}$ has shape $e$; $P_{f}=f$ for all $f\in F$; and $\bigcup_{e\in E(L)}V^{0}(P_{e})=V\setminus V(L)$ (given a path $P$, $V^{0}(P)$ denotes the set of internal vertices of $P$). (Note that the last equality implies that the isolated vertices of $L$ remain isolated in $\mathcal{H}$.) A spanning configuration of shape $(L,F)$ for a multigraph is defined the same way as above. We will need the following result by Girão, Granet, Kühn, Lo, and Osthus.
Lemma 2.10 ([11, Lemma 7.3]).
Let $0<\frac{1}{n}\ll\varepsilon\ll\nu\ll\tau\ll\gamma\ll\eta,\delta\leq 1$. Suppose $\ell\in\mathbb{N}$ satisfies $\ell\leq(\delta-\eta)n$. If $\ell\leq\varepsilon^{2}n$, then let $p\leq n^{-1}$; otherwise, let $p\leq n^{-2}$. Let $D$ and $\Gamma$ be edge-disjoint digraphs on a common vertex set $V$ of size $n$. Suppose that $D$ is $(\delta,\varepsilon)$-almost regular and $\Gamma$ is $(\gamma,\varepsilon)$-almost regular. Suppose further that $\Gamma$ is an $(\varepsilon,p)$-robust $(\nu,\tau)$-outexpander. Let $\mathcal{F}$ be a multiset of directed edges on $V$. Any edge in $\mathcal{F}$ is considered to be distinct from the edges of $D\cup\Gamma$, even if the starting and ending points are the same. Let $F_{1},\dots,F_{\ell}$ be a partition of $\mathcal{F}$. Assume that $(L_{1},F_{1}),\dots,(L_{\ell},F_{\ell})$ are layouts such that $V(L_{i})\subseteq V$ for each $i\in[1,\ell]$ and the following hold, where $L:=\bigcup_{i\in[1,\ell]}L_{i}$.
  1. (a)
    For each $i\in[1,\ell]$, $|V(L_{i})|\leq\varepsilon^{2}n$ and $|E(L_{i})|\leq\varepsilon^{4}n$.
  2. (b)
    Moreover, for each $v\in V$, $d_{L}(v)\leq\varepsilon^{3}n$ and there exist at most $\varepsilon^{2}n$ indices $i\in[1,\ell]$ such that $v\in V(L_{i})$.
Then there exist edge-disjoint submultidigraphs $\mathcal{H}_{1},\dots,\mathcal{H}_{\ell}\subseteq D\cup\Gamma\cup\mathcal{F}$ such that, for each $\in[1,\ell], i\mathcal{H}_{i}$ is a spanning configuration of shape $(L_{i},F_{i})$ and that $D\cup\Gamma-\bigcup_{i\in[1,\ell]}E(\mathcal{H}_{i})$ is a robust $(\nu/2,\tau)$-outexpander.
Results on robust outexpanders can be modified for robust expanders. We will apply the result below and adapt Lemma 2.10 for robust expanders.
Lemma 2.11 ([17, Lemma 3.1]).
Suppose that $0<1/n\ll\eta^{\prime}\ll\nu,\tau,\alpha<1$, and that $G$ is a robust $(\nu,\tau)$-expander on $n$ vertices with $\delta(G)\geq\alpha n$. Then one can orient the edges of $G$ in such a way that the oriented digraph $D$ thus obtained from $G$ satisfies the following properties:
  • (i)
    $D$ is a robust $(\nu/4,\tau)$-outexpander.
  • (ii)
    $d^{+}_{D}(x)=(1\pm\eta^{\prime})d_{G}(x)/2$ and $d^{-}_{D}(x)=(1\pm\eta^{\prime})d_{G}(x)/2$ for every vertex $x$ of $D$.
We now adapt Lemma 2.10 for robust expanders.
Theorem 2.12.
Let $0<\frac{1}{n}\ll\varepsilon^{*}\leq\varepsilon\ll\nu^{\prime}\ll\nu\ll\tau\ll\gamma\ll\alpha,\eta^{*}\ll 1$, and let $\ell\leq\frac{1}{2}(\alpha-\eta^{*})n$. Suppose $G$ is an $(\alpha,\varepsilon^{*})$-almost regular robust $(\nu,\tau)$-expander on $n$ vertices. Suppose that, for each $i\in[1,\ell]$, $F_{i}$ is a multiset of edges on $V(G)$. Assume that $(L_{1},F_{1}),\dots,(L_{\ell},F_{\ell})$ are layouts such that $V(L_{i})\subseteq V(G)$ for each $i\in[1,\ell]$ and the following hold, where $L:=\bigcup_{i\in[1,\ell]}L_{i}$.
  1. (a)
    For each $i\in[1,\ell]$, $|V(L_{i})|\leq\varepsilon^{2}n$ and $|E(L_{i})|\leq\varepsilon^{4}n$.
  2. (b)
    Moreover, for each $v\in V(G)$, $d_{L}(v)\leq\varepsilon^{3}n$ and there exist at most $\varepsilon^{2}n$ indices $i\in[1,\ell]$ such that $v\in V(L_{i})$.
Then there exist edge-disjoint submultigraphs $\mathcal{H}_{1},\ldots,\mathcal{H}_{\ell}\subseteq G+\bigcup_{i\in[1,\ell]}F_{i}$ such that each $\mathcal{H}_{i}$ is a spanning configuration of shape $(L_{i},F_{i})$ and that $G-\bigcup_{i=1}^{\ell}E(\mathcal{H}_{i})$ is a robust $(\nu^{\prime}/2,4\tau)$-expander.
Proof.
Let $\eta^{\prime}$ be defined as in Lemma 2.11 such that $1/n\ll\eta^{\prime}\ll\varepsilon^{*}$. By Lemma 2.11, we can orient the edges of $G$ to get a digraph $D$ such that $D$ is a robust $(\nu/4,\tau)$-outexpander with $d^{+}_{D}(x)=(1\pm\eta^{\prime})d_{G}(x)/2$ and $d^{-}_{D}(x)=(1\pm\eta^{\prime})d_{G}(x)/2$ for every vertex $x$ of $D$. Since $G$ is $(\alpha,\varepsilon^{*})$-almost regular, it follows that $d^{+}_{D}(x),d^{-}_{D}(x)=(\frac{1}{2}\alpha\pm\frac{1}{2}(\varepsilon^{*}+\eta^{\prime}))n$. Thus $D$ is $(\frac{1}{2}\alpha,\frac{1}{2}(\varepsilon^{*}+\eta^{\prime}))$-almost regular. As $\varepsilon^{*}\leq\varepsilon$, $D$ is also $(\frac{1}{2}\alpha,\varepsilon)$-almost regular.
Taking $\lambda=\frac{2\gamma}{\alpha}$ and applying Lemma 2.9, we split $D$ into two edge-disjoint spanning subdigraphs $\Gamma$ and $D^{\prime}$ such that $\Gamma$ is a $(\gamma,\varepsilon)$-almost regular robust $(\gamma\nu/(4\alpha),\tau)$-outexpander, and $D^{\prime}$ is a $(\alpha/2-\gamma,\varepsilon)$-almost regular robust $((\alpha-2\gamma)\nu/(8\alpha),\tau)$-outexpander. As $\nu^{\prime}\ll\nu\ll\gamma\ll\alpha$ and so $\nu^{\prime}$ can be chosen so that $\nu^{\prime}\ll\gamma\nu/(4\alpha)$, by Lemma 2.8, $\Gamma$ is an $(\varepsilon,n^{-2})$-robust $(\nu^{\prime},4\tau)$-outexpander.
Applying Lemma 2.10 with $D^{\prime}$, $\alpha/2-\gamma$, $\nu^{\prime}$, $\varepsilon^{1/5}$, and $\eta^{*}/2$ playing the roles of $D,\delta,\nu,\varepsilon$, and $\eta$, we find edge-disjoint submultidigraphs $\mathcal{H}_{1},\dots,\mathcal{H}_{\ell}\subseteq D^{\prime}\cup\Gamma+\bigcup_{i\in[1,\ell]}F_{i}$ such that, for each $\in[1,\ell], i\mathcal{H}_{i}$ is a spanning configuration of shape $(L_{i},F_{i})$ and $D^{\prime}\cup\Gamma-\bigcup_{i=1}^{\ell}E(\mathcal{H}_{i})$ is a robust $(\nu^{\prime}/2,4\tau)$-outexpander. The latter implies that $G-\bigcup_{i=1}^{\ell}E(\mathcal{H}_{i})$ is a robust $(\nu^{\prime}/2,4\tau)$-expander. ∎
3 Proof of Theorem 1.5
For a graph $G$ and an integer $i\geq 0$, define $V_{i}(G)=\{v\in V(G)\mid d_{G}(v)=i\}$, $W(G)=V(G)\setminus(V_{\delta}(G)\cup V_{\Delta}(G))$, and $g(G)=\Delta(G)-\delta(G)$. We will reduce Theorem 1.5 into the statement below.
Theorem 3.1.
Let $n\in\mathbb{N}$ and suppose that $0<\frac{1}{n}\ll\eta\ll\nu\ll\tau\ll\alpha\ll 1$. Let $G$ be a robust $(\nu,\tau)$-expander on $n$ vertices with $\delta(G)\geq\alpha n$ and $U(G)=\{v\in V(G)\mid\Delta(G)-d_{G}(v)\geq\eta n\}$. Then $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$ if $G$ satisfies one of the following three conditions.
  1. (1)
    $|U(G)|\geq\eta n$ or ($g(G)\geq\eta n$ and $|V_{\delta}(G)\cup V_{\delta+1}(G)|\geq\eta n$);
  2. (2)
    ($|U(G)|<\eta n$, $g(G)\leq 2$, and $|W(G)|\geq 2$) or ($1<g(G)<\eta n$ and $|V_{\delta}(G)\cup V_{\delta+1}(G)|\geq\eta n$);
  3. (3)
    $|U(G)|<\eta n$ and $|W(G)|\leq 1$.
Proof of Theorem 1.5 Assuming Theorem 3.1.
Let $\tau^{*}=\tau(\alpha)$ be defined as in Theorem 1.5. We choose an additional parameter $\nu^{\prime}$ such that
$0<1/n\leq 1/n_{0}\ll\eta\ll\nu^{\prime}\ll\nu\leq\tau\ll\tau^{*}\ll\alpha<1.$
If $G$ meets any of the conditions of Theorem 3.1, then we get $\operatorname{la}(G)\leq\lceil\frac{\Delta(G)+1}{2}\rceil$. Thus we suppose that $|U(G)|<\eta n$, $g(G)\geq 3$, $|W(G)|\geq 2$, and $|V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$. We first show that $|U(G)\cup V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$.
Claim 3.2.
We have $|U(G)\cup V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$.
Proof.
We have $|U(G)|<\eta n$ and $|V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$ by the assumption on $G$. As $V_{\delta}(G)\cup V_{\delta+1}(G)\subseteq U(G)$ if $g(G)\geq\eta n+1$, and $U(G)\subseteq V_{\delta}(G)\cup V_{\delta+1}(G)$ if $g(G)<\eta n+1$, we have $|U(G)\cup V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$ if $g(G)\geq\eta n$. If $g(G)<\eta n$, then we have $U(G)=\emptyset$ and so $|U(G)\cup V_{\delta}(G)\cup V_{\delta+1}(G)|=|V_{\delta}(G)\cup V_{\delta+1}(G)|<\eta n$. ∎
We will remove edge-disjoint paths that each contain all vertices of $V_{\Delta}(G)$ as internal vertices and reduce $G$ to a graph $G^{\prime}$ such that $\delta(G^{\prime})=\delta(G)$ and $G^{\prime}$ satisfies one of the three conditions of Theorem 3.1. Our goal is to apply Theorem 2.12. Thus, we start with identifying edges to be included in $F_{i}$ as needed in the theorem. Let
$W_{0}$ $=$ $W(G),\quad U_{0}=U(G),\quad V_{0}=V_{\delta}(G),$
$Z_{0}$ $=$ $V_{\delta+1}(G),\quad G_{0}^{*}=G-(U_{0}\cup V_{0}\cup Z_{0}),$
$g_{0}$ $=$ $g(G),\quad\text{and}\quad d_{0}(v)=d_{G}(v)\quad\text{for each $v\in V(G)$.}$
We perform the following algorithm. For each $i\geq 1$, if $g_{i-1}\geq 3$, $|W_{i-1}|\geq 2$, and $|U_{i-1}\cup V_{i-1}\cup Z_{i-1}|<\eta n$, then we let
  1. (a)
    $x_{i},y_{i}\in W_{i-1}$ be distinct;
  2. (b)
    $d_{i}(v)=d_{i-1}(v)-2$ for each $v\in V(G_{i-1}^{*})\setminus(V_{i-1}\cup Z_{i-1})$ with $v\not\in\{x_{i},y_{i}\}$, $d_{i}(v)=d_{i-1}(v)-1$ for $v\in\{x_{i},y_{i}\}$, and $d_{i}(v)=d_{G}(v)$ for all other vertices $v$;
  3. (c)
    $g_{i}=g_{i-1}-2$ and $W_{i}=\{v\in V(G)\mid\delta(G)<d_{i}(v)<\Delta(G)-2i\}$;
  4. (d)
    $U_{i}=\{v\in V(G)\mid\Delta(G)-2i-d_{i}(v)\geq\eta n\}$;
  5. (e)
    $V_{i}=\{v\in V(G)\mid d_{i}(v)=\delta(G)\}$ and $Z_{i}=\{v\in V(G)\mid d_{i}(v)=\delta(G)+1\}$;
  6. (f)
    $G_{i}^{*}=G-(U_{i}\cup V_{i}\cup Z_{i})$.
Note that $U_{i}\subseteq U_{0}$ and $V_{0}\cup Z_{0}\subseteq V_{i}\cup Z_{i}$ for each $i$. Although the graph $G_{0}^{*}$ does not contain any vertex of $V_{0}\cup Z_{0}$ initially, it can contain vertices of $V_{i}\cup Z_{i}$ in some later steps. When $g_{0}\geq\eta n$, it is possible that $V(G_{0}^{*})\subseteq V(G_{i}^{*})$ for some $i$; also in the process, it is possible that $V(G_{i}^{*})\subseteq V(G_{0}^{*})$ when $V(G_{0}^{*})\cap(V_{i}\cup Z_{i})\neq\emptyset$.
Since $g_{i}$ is strictly decreasing with $i$, the algorithm above will end after less than $\lfloor g(G)/2\rfloor$ steps. Suppose that upon the completion of the algorithm, all the vertex pairs we found are $x_{1},y_{1},\ldots,x_{\ell},y_{\ell}$. As the process cannot proceed any further, we have $g_{\ell}\leq 2$ or $|W_{\ell}|\leq 1$ or $|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$. The graphs $G_{i}^{*}$ and $G_{j}^{*}$ might be distinct for distinct $i$ and $j$. In order to apply Theorem 2.12 on one single host graph, we match vertices of each $G_{i}^{*}$ that are not in $G_{0}^{*}$ for $i\geq 1$ to some vertices of $G_{0}^{*}$.
For each $i\in[1,\ell]$, we select the following vertices from $G_{0}^{*}$ and also construct $F_{i}$ and $L_{i}$:
  1. (i)
    If $x_{i}\in V(G_{0}^{*})$, let $x_{i}^{\prime}=x_{i}$; otherwise, let $x_{i}^{\prime}x_{i}\in E(G)$ with $x_{i}^{\prime}\in V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$ such that $x_{i}^{\prime}x_{i}$ has not been used in previous steps (this is possible, since $|V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})|\geq(1-2\eta)n$ and $\delta(G)\geq\alpha n$);
  2. (ii)
    If $y_{i}\in V(G_{0}^{*})$, let $y_{i}^{\prime}=y_{i}$; otherwise, let $y_{i}^{\prime}y_{i}\in E(G)$ with $y_{i}^{\prime}\in V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$ such that $y_{i}^{\prime}y_{i}$ has not been used in previous steps and $y^{\prime}_{i}\neq x_{i}^{\prime}$;
  3. (iii)
    We match each vertex $z$ of $V(G_{i}^{*})\setminus V(G^{*}_{0})$ to two distinct vertices $z_{i1},z_{i2}$ of $V(G_{0}^{*})\setminus(\{x_{i}^{\prime},y_{i}^{\prime}\}\cup V_{i-1}\cup Z_{i-1})$ such that $z$ is adjacent in $G$ to both $z_{i1},z_{i2}$, $z_{i1}$ and $z_{i2}$ have not been matched to any other vertex of $V(G_{i}^{*})\setminus V(G^{*}_{0})$, and that $zz_{i1},zz_{i2}$ have not been used in any previous steps (since $V_{0}\cup Z_{0}\subseteq V_{i}\cup Z_{i}$ and $U_{i}\subseteq U_{0}$ for each $i$, we know that $|V(G_{i}^{*})\setminus V(G^{*}_{0})|\leq|U_{0}\setminus U_{i}|<\eta n-1$. As $\delta(G)\geq\alpha n$, vertices $z_{i1},z_{i2}$ exist; also note that when $|V(G_{i}^{*})\setminus V(G^{*}_{0})|>0$ happens, then we have $g_{0}\geq\eta n$);
    In the selection of vertices in (i)-(iii) above, we will make sure that each vertex $v\in V(G_{0}^{*})$ is used in at most $\sqrt{\eta}n$ steps. This is possible: for each $F_{i}$, we select at most $2+2(\eta n-1)=2\eta n$ distinct vertices of $V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$; there are at most $3\sqrt{\eta}\ell\leq 3\sqrt{\eta}(1-\alpha)n/2$ vertices of $V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$ that have been used in at least $\lfloor\sqrt{\eta}n\rfloor$ previous steps; and each vertex from $\{x_{i},y_{i},z\}$ has in $G$ at least $(\alpha-2\eta)n$ distinct neighbors from $V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$. Since $(\alpha-2\eta)n-3\sqrt{\eta}(1-\alpha)n/2>2\eta n$, we can find the desired vertices at Step $i$ from $V(G_{0}^{*})\setminus(V_{i-1}\cup Z_{i-1})$ that are used at most $\lfloor\sqrt{\eta}n\rfloor-1$ times in the previous steps.
  4. (iv)
    Let $F_{i}$ be the linear forest consisting of edges $x_{i}^{\prime}y_{i}^{\prime}$ and $z_{i1}z_{i2}$ for each $z\in V(G_{i}^{*})\setminus V(G^{*}_{0})$. Since $|V(G_{i}^{*})\setminus V(G^{*}_{0})|\leq|U_{0}\setminus U_{i}|<\eta n-1$, $e(F_{i})<\eta n$ for each $i\in[1,\ell]$. Furthermore, by the comment on the selection of vertices before (iv), we know that every vertex of $G_{0}^{*}$ is used by at most $\sqrt{\eta}n$ distinct $F_{i}$’s.
  5. (v)
    For each $i\in[1,\ell]$, let $v_{i1},v_{i2}\in V(G_{0}^{*})\setminus(V(F_{i})\cup V_{i-1}\cup Z_{i-1})$ be distinct and such that, for any $v\in V(G_{0}^{*})$, there exist at most two $(i,j)\in[1,\ell]\times[1,2]$ such that $v=v_{ij}$. For each $i\in[1,\ell]$, denote by $P_{i,1},\dots,P_{i,f_{i}}$ the components of $F_{i}$ and, for $j\in[1,f_{i}]$, denote by $u_{ij}$ and $w_{ij}$ the starting and ending vertices of $P_{i,j}$. For each $i\in[1,\ell]$, let $L_{i}$ be the union of
    $\{v_{i1}u_{i1}P_{i,1}w_{i1}u_{i2}P_{i,2}w_{i2}u_{i3}\ldots w_{i,f_{i}}v_{i2},v_{i2}v_{i1}\},$
    and $\left((V_{i-1}\cup Z_{i-1})\cap V(G_{0}^{*})\right)\setminus\{x_{i},y_{i}\}$ (those vertices are isolated vertices in $L_{i}$, and their existence implies $g_{i-1}\leq\eta n+1$).
    Denote $L=\bigcup_{i\in[1,\ell]}L_{i}$. Note that $(L_{1},F_{1}),\dots,(L_{\ell},F_{\ell})$ are layouts such that, for each $i\in[1,\ell]$, $V(L_{i})\subseteq V(G_{0}^{*})$, $|V(L_{i})|\leq 2e(F_{i})+2+|V_{i-1}\cup Z_{i-1}|\leq 3\eta n$ and $e(L_{i})\leq 2e(F_{i})+3<3\eta n$. Moreover, for each $v\in V(G_{0}^{*})$, $d_{L}(v)\leq 2\sqrt{\eta}n+2$ and there exist at most $\sqrt{\eta}n+2+\eta n+1<2\sqrt{\eta}n$ indices $i\in[1,\ell]$ such that $v\in V(L_{i})$ (when $(V_{i}\cup Z_{i})\cap V(G_{0}^{*})\neq\emptyset$, we have $g_{i}\leq\eta n+1$. Thus a vertex in $V_{i}\cup Z_{i}$ can be used by at most $\eta n$ other $L_{j}$’s for $j\geq i$).
Let $\alpha^{*}=\Delta(G)/n$. As $d_{G}(v)\geq(\alpha^{*}-\eta)n$ for any $v\in V(G_{0}^{*})$, by Claim 3.2, $G_{0}^{*}$ is $(\alpha^{*},2\eta)$-almost regular. Furthermore, $G_{0}^{*}$ is a robust $(\nu-\eta,2\tau)$-expander by Lemma 2.6. Applying Theorem 2.12 with $G_{0}^{*}$, $\alpha^{*}$, $\nu-2\eta$, $2\tau$, and $\alpha$ playing the roles of $G,\alpha,\nu,\tau$, and $\eta^{*}$, and with $\varepsilon^{*}=2\eta$ and $\varepsilon=\eta^{1/7}$, we find edge-disjoint subgraphs $C_{1},\ldots,C_{\ell}$ of $G_{0}^{*}\cup(\bigcup_{i=1}^{\ell}F_{i})$ such that $C_{i}$ is a spanning configuration of shape $(L_{i},F_{i})$ and $E(F_{i})\subseteq E(C_{i})$.
For each $i\in[1,\ell]$, for $x_{i}^{\prime}y_{i}^{\prime}\in E(F_{i})\cap E(C_{i})$, we delete $x_{i}^{\prime}y_{i}^{\prime}$ from $C_{i}$, and add the edges $x_{i}^{\prime}x_{i}$ (if $x_{i}\neq x_{i}^{\prime}$) and $y_{i}^{\prime}y_{i}$ (if $y_{i}\neq y_{i}^{\prime}$) to the resulting path; for any other edge $z_{i1}z_{i2}$ of $F_{i}$, we let $z\in V(G_{i}^{*})\setminus V(G^{*}_{0})$ such that $z$ was matched to $z_{i1},z_{i2}$ in step (iii) above, then we replace $z_{i1}z_{i2}$ on $C_{i}$ by the path $z_{i1}zz_{i2}$. Denote by $P_{i}$ the resulting path. Then $P_{i}$ is a Hamilton $(x_{i},y_{i})$-path of $G[V(G_{i}^{*})\cup\{x_{i},y_{i}\}]$. Let $G^{\prime}=G-\bigcup_{i=1}^{\ell}E(P_{i})$. As $G^{*}_{0}-\bigcup_{i=1}^{\ell}E(C_{i})$ is a robust $(\nu^{\prime}/2,8\tau)$-expander by Theorem 2.12, $G^{\prime}$ is a robust $(\nu^{\prime}/2-\eta,16\tau)$-expander by Lemma 2.6. As $\tau\ll\tau^{*}$, $G^{\prime}$ is also a robust $(\nu^{\prime}/2-\eta,\tau^{*})$-expander. Note that $g(G^{\prime})=g_{\ell}$, $W(G^{\prime})=W_{\ell}$, $U(G^{\prime})=U_{\ell},V_{\delta}(G^{\prime})=V_{\ell}$, $V_{\delta+1}(G^{\prime})=Z_{\ell}$, $U(G^{\prime})\subseteq U(G)$, and $\delta(G^{\prime})=\delta(G)\geq\alpha n$. Since $|U(G^{\prime})|\leq|U(G)|<\eta n$, and ($g_{\ell}\leq 2$ or $|W_{\ell}|\leq 1$ or $|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$), we have one of the following three possibilities:
  1. (A)
    $|W_{\ell}|\leq 1$. In this case, Theorem 3.1(3) holds.
  2. (B)
    $|W_{\ell}|\geq 2$ and $g_{\ell}\leq 2$. In this case, Theorem 3.1(2) holds.
  3. (C)
    $|W_{\ell}|\geq 2$, $g_{\ell}\geq 3$, and $|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$.
    • If $g_{\ell}\geq\eta n+1$, then we have $V_{\ell}\cup Z_{\ell}\subseteq U_{\ell}$ and so $|U_{\ell}|=|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$. This gives a contradiction to $|U_{\ell}|\leq|U(G)|<\eta n$.
    • If $\eta n\leq g_{\ell}<\eta n+1$, then we have $U_{\ell}=V_{\ell}$ and so $|V_{\ell}\cup Z_{\ell}|=|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$. In this case, Theorem 3.1(1) holds.
    • If $g_{\ell}<\eta n$, then $U_{\ell}=\emptyset$ and so $|V_{\ell}\cup Z_{\ell}|=|U_{\ell}\cup V_{\ell}\cup Z_{\ell}|\geq\eta n$. Since also $g_{\ell}\geq 3$, in this case, Theorem 3.1(2) holds.
Applying Theorem 3.1, we get $\operatorname{la}(G^{\prime})\leq\lceil\frac{1}{2}(\Delta(G^{\prime})+1)\rceil$. Thus $\operatorname{la}(G)\leq\ell+\lceil\frac{1}{2}(\Delta(G^{\prime})+1)\rceil=\lceil\frac{1}{2}(\Delta(G)+1)\rceil$. ∎
The following lemma will be used in the proof of Theorem 3.1 when the second condition is met. For a positive integer $d$, the $d$-deficiency of a vertex $v$ in $G$ is $\operatorname{def}_{G}(v,d)=\max\{d-d_{G}(v),0\}$.
Lemma 3.3.
Let $n\in\mathbb{N}$ and suppose that $0<1/n\ll\nu\leq\tau\ll\alpha<1$. Suppose that $G$ is a robust $(\nu,\tau)$-expander on $n$ vertices with $\delta(G)\geq\alpha n$, and that for some positive integer $d$, $\operatorname{def}_{G}(v_{1},d)\geq\ldots\geq\operatorname{def}_{G}(v_{n},d)$, and $\operatorname{def}_{G}(v_{1},d)\leq\nu^{6}n+1$. If $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},d)$ is even and $\operatorname{def}_{G}(v_{1},d)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},d)$, then there exists some integer $\ell\leq 3\nu^{2}n$ such that after removing $\ell$ edge-disjoint linear forests from $G$, we can get a graph $G^{*}$ such that $d_{G^{*}}(v_{i})=d_{G}(v_{i})-2\ell+\operatorname{def}_{G}(v_{i},d)$ for each $i\in[1,n]$, and $G^{*}$ is a robust $(\nu/2,\tau)$-expander with $\delta(G^{*})\geq(\alpha-6\nu^{2})n$.
Proof.
By Theorem 2.1, there exists a multigraph $H$ on $V(G)$ such that $d_{H}(x_{i})=\operatorname{def}_{G}(v_{i},d)$ for each $i\in[1,n]$. The multigraph $H$ will aid us to find a desired subgraph $G^{*}$ of $G$. Note that $\Delta(H)=\operatorname{def}_{G}(v_{1},d)=d-d_{G}(v_{1})\leq\nu^{6}n+1$.
By Vizing’s Theorem on chromatic index, we have $\chi^{\prime}(H)\leq\Delta(H)+\mu(H)\leq 2\Delta(H)\leq 2\nu^{6}n+2$, where $\mu(H)$ is the maximum number of edges joining two vertices in $H$. By further partitioning the edges in each color class of a $(2\nu^{6}n+2)$-edge coloring of $H$ into $\nu^{-4}$ matchings, we can greedily partition the edges of $H$ into
$\ell\leq\frac{2\nu^{6}n+2}{\nu^{4}}<3\nu^{2}n$
matchings $M_{1},M_{2},\ldots,M_{\ell}$ each of size at most
$\frac{\frac{n}{2}}{\nu^{-4}}$ $=$ $\nu^{4}n/2.$
Now we take out linear forests from $G$ by applying Lemma 2.7 with $M_{1},M_{2},\ldots,M_{\ell}$ iteratively. More precisely, we define spanning subgraphs $G_{0},G_{1},\ldots,G_{\ell}$ of $G$ and edge-disjoint linear forests $F_{1},F_{2},\ldots,F_{\ell}$ such that
  1. (a)
    $G_{0}:=G$ and $G_{i}:=G_{i-1}-E(F_{i})$ for $i\geq 1$.
  2. (b)
    $F_{i}$ is a spanning linear forest in $G_{i-1}$ whose leaves are precisely the vertices in $V(M_{i})$.
Let $G_{0}=G$ and suppose that for some $i\in[1,\ell]$, we have already defined $G_{0},G_{1},\ldots,G_{i-1}$ and $F_{1},\ldots,F_{i-1}$. Since $\Delta(F_{1}\cup\ldots\cup F_{i-1})\leq 2(i-1)\leq 2(\ell-1)\leq 6\nu^{2}n$, it follows that $G_{i-1}$ is still a robust $(\nu-6\nu^{2},\tau)$-expander and so is a robust $(2^{-1/4}\nu,\tau)$-expander. Since $M_{i}$ has size at most $\nu^{4}n/2$, we can apply Lemma 2.7 to $G_{i-1}$ and $M_{i}$ to obtain a spanning linear forest $F_{i}$ in $G_{i-1}$ whose leaves are precisely the vertices in $V(M_{i})$. Let $G_{i}=G_{i-1}-E(F_{i})$.
We claim that $G_{\ell}$ satisfies the degree constraint for Lemma 3.3. Consider any vertex $u\in V(G_{\ell})$. For every $i\in[1,\ell]$, $d_{F_{i}}(u)=1$ if $u$ is an endvertex of some edge of $M_{i}$ and $d_{F_{i}}(u)=2$ otherwise. Since $M_{1},M_{2},\ldots,M_{\ell}$ partition $E(H)$, we know that $\sum\limits_{i=1}^{\ell}d_{F_{i}}(u)=2\ell-d_{H}(u)=2\ell-\operatorname{def}_{G}(u,d)$. Thus
$d_{G_{\ell}}(u)=d_{G}(u)-\sum\limits_{i=1}^{\ell}d_{F_{i}}(u)=d_{G}(u)-2\ell+\operatorname{def}_{G}(u,d).$
Furthermore $\delta(G_{\ell})\geq(\alpha-6\nu^{2})n$. Thus letting $G^{*}=G_{\ell}$ gives us the desired subgraph. ∎
Proof of Theorem 3.1.
Let $\tau^{*}=\max\{\tau(\alpha/2),\tau(\alpha),\tau(\alpha-6\nu^{2})\}$, where $\tau(\cdot)$ is defined in Theorem 1.2. We choose an additional parameter $\nu^{\prime}$ such that
$0<1/n\leq 1/n_{0}\ll\eta\ll\nu^{\prime}\ll\nu\leq\tau\ll\tau^{*}\ll\alpha<1,$
where $n_{0}\geq\max\{N_{0}(\alpha/2,\eta^{2},\tau^{*}),N_{0}(\alpha-6\nu^{2},\nu^{\prime}/3,\tau^{*}),N_{0}(\alpha,\nu/2,\tau^{*})\}$ satisfying $0<1/n_{0}\ll\eta$ and $N_{0}(\cdot)$ is defined in Theorem 1.2.
Let $G$ be a robust $(\nu,\tau)$-expander on $n\geq n_{0}$ vertices with $\delta(G)\geq\alpha n$, and let
$\Delta=\Delta(G),\quad\delta=\delta(G),\quad V_{\delta}=V_{\delta}(G),\quad V_{\delta+1}=V_{\delta+1}(G),\quad W=W(G),\quad g=g(G).$
Additionally, define
$U=\{v\in V(G)\mid\Delta(G)-d_{G}(v)\geq\eta n\}.$
For any two vertices of degree less than $\Delta$, if they are not adjacent in $G$, we add an edge joining them to $G$. By iterating this process, we thus assume that all vertices with degrees less than $\Delta$ are mutually adjacent in $G$. Thus, by Lemma 2.3, we have $|V_{\Delta}|>\frac{n}{2}$. Since $W\cup V_{\delta}$ is a clique in $G$, we have $\delta\geq|W\cup V_{\delta}|-1$. Thus
$|V_{\Delta}|>\max\{n/2,n-\delta-2\}.$ (1)
If $G$ is regular, then we are done by Theorem 1.3. Thus we assume that $G$ is not regular and so $g\geq 1$. This together with the fact that $W\cup V_{\delta}$ is a clique in $G$ imply
$\Delta\leq n-2.$ (2)
We let $V(G)=\{v_{1},\ldots,v_{n}\}$, and we assume, without loss of generality, that
$d_{G}(v_{1})\leq\ldots\leq d_{G}(v_{n}).$
We proceed with the proof according to the three given cases.
Case 1: $|U|\geq\eta n$ or ($g\geq\eta n$ and $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$).
If $\Delta$ is odd, we delete a perfect matching $M$ from $G$ if $n$ is even, and we delete a matching covering all vertices of $V(G)\setminus\{v_{1}\}$ if $n$ is odd, where recall that we assumed $d_{G}(v_{1})=\delta$ and $M$ exists as $G$ is Hamiltonian by Lemma 2.7 with $k=1$. Thus in the following, we let $\Delta^{*}=\Delta$ and $G^{*}=G$ if $\Delta$ is even, and $\Delta^{*}=\Delta-1$ and $G^{*}=G-M$ if $\Delta$ is odd. Note that $\sum_{i=1}^{n}\operatorname{def}_{G^{*}}(v_{i},\Delta^{*})\geq(\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta))-1$, here we have this extra “1-1” in the bound since when both $\Delta$ and $n$ are odd, we have $\Delta^{*}-d_{G^{*}}(v_{1})=\Delta-d_{G}(v_{1})-1$. We will add vertices to $G^{*}$ and construct a simple $\Delta^{*}$-regular graph $H$ in three steps such that $G^{*}\subseteq H$, and $H$ is a robust $(\eta^{2},2\tau)$-expander.
Let $\operatorname{def}(G^{*})=\sum_{i=1}^{n}\operatorname{def}_{G^{*}}(v_{i},\Delta^{*})$. Then by the condition of this case and the way of reducing $G$ to $G^{*}$, we have $\operatorname{def}(G^{*})\geq\eta n(\eta n-1)-1\geq\eta^{2}n^{2}/2\geq\Delta^{*}+5$. We add a set $X$ of new vertices to $G^{*}$ such that $|X|\geq\frac{\Delta^{*}+4}{2}$ is the smallest integer satisfying the following properties:
  1. (i)
    $|X|\equiv n\ (\mbox{mod}\ 2)$;
  2. (ii)
    $|X|\geq\Delta^{*}-\lfloor\frac{{\rm def}(G^{*})}{|X|}\rfloor+4$.
Note that when $|X|=\Delta^{*}+4$ or $|X|=\Delta^{*}+5$, the conditions above are satisfied. Thus a smallest integer $|X|$ satisfying the conditions above exists and $|X|\leq\Delta^{*}+5\leq n+4$.
Let $\delta^{*}=\delta(G^{*})$. We claim that $(\Delta^{*}+4)^{2}-4\operatorname{def}(G^{*})>(\Delta^{*}-2\delta^{*})^{2}$. To see this, we have $\operatorname{def}(G^{*})\leq(\Delta^{*}-\delta^{*})\times|W\cup V_{\delta}|<(\Delta^{*}-\delta^{*})(\delta^{*}+2)$ by $|V_{\Delta}|>\max\{n/2,n-\delta-2\}$ from (1) (the set of maximum degree vertices of $G$ and $G^{*}$ stay the same) and so
$(\Delta^{*}+4)^{2}-4\operatorname{def}(G^{*})$ $\geq$ $(\Delta^{*}+4)^{2}-4\Delta^{*}\delta^{*}+4(\delta^{*})^{2}-8(\Delta^{*}-\delta^{*})$ (3)
$=$ $(\Delta^{*}-2\delta^{*})^{2}+8\delta^{*}+16>0.$
Now from (ii) that $|X|\geq\Delta^{*}-\lfloor\frac{{\rm def}(G^{*})}{|X|}\rfloor+4$, we get $|X|^{2}-(\Delta^{*}+4)|X|+\operatorname{def}(G^{*})\geq 0$. As $|X|\geq\frac{\Delta^{*}+4}{2}$ and (3), solving the quadratic inequality, we get
$|X|$ $\geq$ $\frac{\Delta^{*}+4+\sqrt{(\Delta^{*}+4)^{2}-4\operatorname{def}(G^{*})}}{2}$ (4)
$>$ $\frac{\Delta^{*}+4+\sqrt{(\Delta^{*}-2\delta^{*})^{2}}}{2}$
$>$ $\Delta^{*}-\delta^{*},$
where the last inequality above holds regardless whether $\Delta^{*}-2\delta^{*}\geq 0$.
Let $d=\Delta^{*}-(\lceil\frac{{\rm def}(G^{*})}{|X|}\rceil-1)$. We also claim that $d-1\geq|X|-6$. If $|X|<\frac{\Delta^{*}}{2}+4$, then as $|X|\geq\frac{\Delta^{*}+4}{2}$ and $\operatorname{def}(G^{*})<\frac{1}{4}(\Delta^{*}+4)^{2}$, we have $d-1=\Delta^{*}-\lceil\frac{{\operatorname{def}}(G^{*})}{|X|}\rceil+1-1\geq\frac{\Delta^{*}}{2}-2>|X|-6$. If $|X|\geq\frac{\Delta^{*}}{2}+4$, then by the minimality requirement on the size of $X$, we know that $|X|-2<\Delta^{*}-\lfloor\frac{{\rm def}(G^{*})}{|X|}\rfloor+4$ and so $|X|-6\leq\Delta^{*}-\lceil\frac{{\operatorname{def}}(G^{*})}{|X|}\rceil+1-1=d-1$.
Suppose that ${\rm def}(G^{*})\equiv\ell\ (\mbox{mod}\ |X|)$, where $\ell\in[0,|X|-1]$. Let $d_{1}=\ldots=d_{\ell}=\Delta^{*}-\lceil\frac{{\rm def}(G^{*})}{|X|}\rceil=d-1$, and $d_{\ell+1}=\ldots=d_{|X|}=\Delta^{*}-(\lceil\frac{{\rm def}(G^{*})}{|X|}\rceil-1)=d$. Since $|X|\equiv n\ (\mbox{mod}\ 2)$, we know that
$\sum\limits_{i=1}^{|X|}d_{i}=\Delta^{*}|X|-{\rm def}(G^{*})\equiv\Delta^{*}n-{\rm def}(G^{*})=2e(G^{*})\equiv 0\ (\mbox{mod}\ 2).$
Furthermore, we have
$d=\Delta^{*}-\left(\left\lceil\frac{{\operatorname{def}}(G^{*})}{|X|}\right\rceil-1\right)\geq\Delta^{*}-\frac{(\Delta^{*}+4)^{2}/4}{(\Delta^{*}+4)/2}=\Delta^{*}-\frac{\Delta^{*}+4}{2}>2.$
Let $X=\{x_{1},\ldots,x_{|X|}\}$. By Lemma 2.2, there is a graph $R$ on $X$ such that $d_{R}(x_{i})=d_{i}$ for each $i\in[1,|X|]$. We now add ${\rm{def}}(G^{*})$ edges joining vertices from $V(G^{*})\setminus V_{\Delta}$ to $X$. We list these ${\rm{def}}(G^{*})$ edges as $e_{1},e_{2},\ldots,e_{{\rm{def}}(G^{*})}$ and add them one by one such that
  1. (a)
    For each vertex $v\in V(G^{*})\setminus V_{\Delta}$, the ${\rm def}_{G^{*}}(v)$ edges joining $v$ with the vertices of $X$ are listed consecutively in the ordering above.
  2. (b)
    For each $x_{i}\in X$, $x_{i}$ is incident with $\Delta^{*}-d_{i}$ edges for each $i\in[1,|X|]$.
The resulting multigraph is denoted by $H$. Then $H$ is $\Delta^{*}$-regular by the above construction. The multigraph $H$ is actually simple, since $|X|>\Delta^{*}-\delta^{*}$ by (4). Let $|V(H)|=n+|X|=n^{\prime}$. By $|X|\leq n+4$ and $\Delta^{*}\geq\delta+\eta n-1>\alpha n+2$, it follows that $\Delta^{*}>\alpha n^{\prime}/2$. In the rest of the proof, we mainly show that $H$ is a robust $(\eta^{2},2\tau)$-expander. Let $S\subseteq V(H)$ with $2\tau n^{\prime}\leq|S|\leq(1-2\tau)n^{\prime}$. We show that $|RN_{H}(S)|\geq|S|+\eta^{2}n^{\prime}$.
Consider first that $|S\cap V(G^{*})|\geq\tau n$. Since $G^{*}$ is a robust $(\nu/2,\tau)$-expander implies it is a $(5\eta^{2},\tau)$-expander, we have $|RN_{G^{*}}(S)|\geq|S\cap V(G^{*})|+5\eta^{2}n$. If $|S\cap X|\leq\eta^{2}n^{\prime}+6$, then
[…]
$\geq|S|+\eta^{2}n^{\prime},$
as $|X|\leq n+4$ and so $n^{\prime}\leq 2n+4$. Thus we assume that $|S\cap X|>\eta^{2}n^{\prime}+6$. Then as $\delta(R)=d-1$ and $d-1\geq|X|-6$, we get $X\subseteq RN_{H}(S)$. Therefore,
$|RN_{H}(S)|\geq|RN_{G^{*}}(S)|+|X\cap S|\geq|S\cap V(G^{*})|+5\eta^{2}n+|X\cap S|\geq|S|+\eta^{2}n^{\prime}.$
Consider then that $|S\cap V(G^{*})|<\tau n$. As $|S|\geq 2\tau n^{\prime}$, we have that $|S\cap X|>\tau n>\eta^{2}n^{\prime}+6$. Again, as $\delta(R)=d-1\geq|X|-6$, we have $X\subseteq RN_{H}(S)$.
Assume for now that $|S\cap V(G^{*})|>\eta n/2$. Then as $\delta(G^{*})\geq\alpha n-1$, we get
$|S\cap V(G^{*})|(\alpha n-1)$ $\leq$ $\sum_{v\in S\cap V(G^{*})}d_{G^{*}}(v)$
$\leq$ $|RN_{G^{*}}(S\cap V(G^{*}))|\times|S\cap V(G^{*})|+\eta^{2}n^{\prime}\times n.$
This gives $|RN_{G^{*}}(S\cap V(G^{*}))|\geq(\alpha n-1)-\frac{\eta^{2}n^{\prime}\times n}{|S\cap V(G^{*})|}\geq(\alpha n-1)-5\eta n>\tau n+\eta^{2}n^{\prime}$. Thus
$|RN_{H}(S)|\geq|X|+|RN_{G^{*}}(S\cap V(G^{*}))|\geq|X\cap S|+\tau n+\eta^{2}n^{\prime}\geq|S|+\eta^{2}n^{\prime}.$
Assume then that $|S\cap V(G^{*})|<\eta n/2$. If $|X\backslash S|\geq\eta^{2}n^{\prime}+\eta n/2$, then
$|RN_{H}(S)|\geq|X|=|S|-|S\cap V(G^{*})|+|X\setminus S|\geq|S|-\eta n/2+\eta^{2}n^{\prime}+\eta n/2=|S|+\eta^{2}n^{\prime}.$
Therefore, we have $|X\backslash S|<\eta^{2}n^{\prime}+\eta n/2$. Let $U^{*}=U$ if $|U|\geq\eta n$ and $U^{*}=V_{\delta}\cup V_{\delta+1}$ otherwise. Since every vertex from $U^{*}$ has in $H$ at least $\eta n-1$ neighbors from $X$, we then know that $U^{*}\subseteq RN_{H}(S)$ and so
$|RN_{H}(S)|\geq|X|+|U^{*}|\geq|X\cap S|+|V(G^{*})\cap S|+\eta n/2\geq|S|+\eta^{2}n^{\prime}.$
Thus $H$ is a $\Delta^{*}$-regular robust $(\eta^{2},2\tau)$-expander, where $\Delta^{*}>\alpha n^{\prime}/2$ is even. As $\tau\ll\tau^{*}$, $H$ is a robust $(\eta^{2},2\tau)$-expander implies that $H$ is a robust $(\eta^{2},\tau^{*})$-expander. As $n$ is taken to be at least $N_{0}(\alpha/2,\eta^{2},\tau^{*})$ as defined in Theorem 1.2, the theorem implies that $H$ has a Hamilton decomposition into $\Delta^{*}/2$ Hamilton cycles. By deleting the edges incident with vertices of $X$ from the decomposition, we obtain a decomposition of $G^{*}$ into $\Delta^{*}/2$ linear forests. If $\Delta=\Delta^{*}$, we get $\operatorname{la}(G)\leq\Delta/2$. Otherwise, we get $\operatorname{la}(G)\leq 1+(\Delta-1)/2=(\Delta+1)/2$.
Case 2: ($|U|<\eta n$, $g\leq 2$, and $|W|\geq 2$) or ($1<g<\eta n$ and $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$).
Since $d_{G}(v_{1})\leq\ldots\leq d_{G}(v_{n})$, we have $\operatorname{def}_{G}(v_{1},\Delta)\geq\ldots\geq\operatorname{def}_{G}(v_{n},\Delta)$. If $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is even and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, applying Lemma 3.3 with $d=\Delta$, we can reduce $G$ into a regular graph $G^{*}$ with degree $\Delta-2\ell$ by taking out $\ell\leq 3\nu^{2}n$ edge-disjoint linear forests. Then $G^{*}$ is still a robust $(\nu/2,\tau)$-expander and so a robust $(\nu/2,\tau^{*})$-expander with $\delta(G^{*})\geq(\alpha-6\nu^{2})n$. By Theorem 1.3, we have $\operatorname{la}(G^{*})\leq\lceil(\Delta-2\ell+1)/2\rceil$. Thus $\operatorname{la}(G)\leq\ell+\lceil(\Delta-2\ell+1)/2\rceil=\lceil(\Delta+1)/2\rceil$.
Thus we assume that $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd or $\operatorname{def}_{G}(v_{1},\Delta)>\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. If $|V_{\delta}\cup V_{\delta+1}|\geq\eta n$, as $g>1$, then we always have $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. If $|V_{\delta}\cup V_{\delta+1}|<\eta n$, $g\leq 2$, and $|W|\geq 2$, then as $\operatorname{def}_{G}(v_{1},\Delta)\leq 2$, and $\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)\geq\sum_{w\in W}\operatorname{def}_{G}(w,\Delta)\geq 2$, it follows that $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. Therefore, $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$. Since $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)=n\Delta-2e(G)$, it follows that both $n$ and $\Delta$ are odd. Since $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)$ is odd and $\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)$, we know $\operatorname{def}_{G}(v_{3},\Delta)>0$. Let $P$ be a Hamilton $(v_{1},v_{n})$-path in $G$ by Lemma 2.7. Let $G_{1}=G-E(P)$. Then we have $d_{G_{1}}(v)=d_{G}(v)-2$ for $v\in V(G_{1})$ with $v\not\in\{v_{1},v_{n}\}$, $d_{G_{1}}(v_{1})=\delta-1$ and $d_{G_{1}}(v_{n})=\Delta-1$. Thus $\Delta(G_{1})=\Delta-1$ and $\delta-2\leq\delta(G_{1})\leq\delta-1$.
For every vertex $v\in V_{\Delta}$, by the definition, we have $\operatorname{def}_{G_{1}}(v,\Delta-2)=\operatorname{def}_{G}(v,\Delta)=0$; for every vertex $v\in V(G)\setminus V_{\Delta}$ with $v\neq v_{1}$, we have $\operatorname{def}_{G_{1}}(v,\Delta-2)=\operatorname{def}_{G}(v,\Delta)$, and $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)=\operatorname{def}_{G}(v_{1},\Delta)-1=\Delta-\delta-1=g-1<\eta n-1$. Therefore, we get $\sum_{i=1}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)=\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta)-1$ and so $\sum_{i=1}^{n}\operatorname{def}_{G}(v_{i},\Delta-2)$ is even. We next show that the largest value among $\operatorname{def}_{G_{1}}(v_{1},\Delta-2),\ldots,\operatorname{def}_{G_{1}}(v_{n},\Delta-2)$ is at most the sum of the rest values. If $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)\geq\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$, then $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)$ is a largest value among $\operatorname{def}_{G_{1}}(v_{1},\Delta-2),\ldots,\operatorname{def}_{G_{1}}(v_{n},\Delta-2)$. Then we have $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)<\operatorname{def}_{G}(v_{1},\Delta)\leq\sum_{i=2}^{n}\operatorname{def}_{G}(v_{i},\Delta)=\sum_{i=2}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)$. Otherwise, we have $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)<\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$. Then we have $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)=\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+1$ as $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+1=\operatorname{def}_{G}(v_{1},\Delta)\geq\operatorname{def}_{G}(v_{2},\Delta)=\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$. In this case, $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)$ is a largest value among $\operatorname{def}_{G_{1}}(v_{1},\Delta-2),\ldots,\operatorname{def}_{G_{1}}(v_{n},\Delta-2)$. Since $\operatorname{def}_{G_{1}}(v_{3},\Delta-2)=\operatorname{def}_{G}(v_{3},\Delta)>0$, we have $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)\leq\operatorname{def}_{G_{1}}(v_{1},\Delta-2)+\sum_{i=3}^{n}\operatorname{def}_{G_{1}}(v_{i},\Delta-2)$. Furthermore, we have $\delta(G_{1})\geq\delta(G)-1\geq\alpha^{*}n$ where $\alpha^{*}\geq 0.99\alpha$, $\operatorname{def}_{G_{1}}(v_{1},\Delta-2)\leq\Delta-\delta+1<\eta n+1$, and $\operatorname{def}_{G_{1}}(v_{2},\Delta-2)<\eta n$. Therefore, the graph $G_{1}$ and sequence of integers $\operatorname{def}_{G_{1}}(v_{1},\Delta-2),\ldots,\operatorname{def}_{G_{1}}(v_{n},\Delta-2)$ satisfy the conditions of Lemma 3.3.
Now applying Lemma 3.3 with $G_{1}$ in the place of $G$ and with $d=\Delta-2$, we can reduce $G_{1}$ into a graph $G^{*}$ by taking out $\ell\leq 3\nu^{2}n$ edge-disjoint linear forests such that for any $v\in V(G_{1})$, we have $d_{G^{*}}(v)=d_{G_{1}}(v)-2\ell+\operatorname{def}_{G_{1}}(v,\Delta-2)$. For any $i\in[1,n-1]$, as $d_{G_{1}}(v_{i})\leq\Delta-2$, it follows that $d_{G_{1}}(v_{i})+\operatorname{def}_{G_{1}}(v_{i},\Delta-2)=\Delta-2$. Thus all vertices except $v_{n}$ of $G^{*}$ have degree $\Delta-2-2\ell$ and $v_{n}$ has degree $\Delta-1-2\ell$, and $G^{*}$ is a robust $(\nu/2,\tau)$-expander.
Applying Lemma 2.5, $\overline{G^{*}}$ has a matching excluding the vertex $v_{n}$ and covering at least $n-\frac{n}{n-(\Delta-2-2\ell)}-3$ vertices of $G^{*}$. Since $\ell\leq 3\nu^{2}n$ and $\Delta\geq\alpha n$, we have $(\Delta-2-2\ell)+1>\frac{n}{n-(\Delta-2-2\ell)}+2$, and so $n-\frac{n}{n-(\Delta-2-2\ell)}-3>n-1-((\Delta-2-2\ell)+1)$. Thus $\overline{G^{*}}$ has a matching $M$ covering exactly $n-1-((\Delta-2-2\ell)+1)$ vertices excluding the vertex $v_{n}$. We let $G_{2}$ be obtained from $G^{*}+M$ by adding a new vertex $x$ and all the edges joining $x$ and vertices from $V(G^{*})\setminus(V(M)\cup\{v_{n}\})$. The graph $G_{2}$ is $((\Delta-2-2\ell)+1)$-regular and is a robust $(\nu/2,\tau)$-expander and so a robust $(\nu/2,\tau^{*})$-expander. Thus $G_{2}$ has a Hamilton decomposition by Theorem 1.2. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G^{*}$ into $\lceil((\Delta-2-2\ell)+1)/2\rceil$ linear forests. Thus $\operatorname{la}(G)\leq 1+\ell+\lceil((\Delta-2-2\ell)+1)/2\rceil=\lceil(\Delta+1)/2\rceil$, as desired.
Case 3: $|U|<\eta n$ and $|W|\leq 1$.
We may assume that $|V_{\delta}|<\eta n$. Indeed, if $g\geq\eta n$, then we have $V_{\delta}\subseteq U$ and so $|V_{\delta}|<\eta n$; and if $g<\eta n$ and $|V_{\delta}|\geq\eta n$, then we are done by Case 2, thus $|V_{\delta}|<\eta n$ again.
As $g=\Delta-\delta\leq n-2-\delta$ by Inequality (2), we have $|V_{\Delta}|\geq g+1$ by Inequality (1). If $|W|=1$, let $w\in W$ and $g_{0}=d_{G}(w)-\delta$. We will remove from $G$ some edge-disjoint paths and reduce it to a graph whose maximum degree and minimum degree differ by exactly one. Recall that by our definition of $W$, we have $V_{\Delta}=V(G)\setminus(V_{\delta}\cup W)$.
We will select $2\lceil g/2\rceil$ vertices $x_{1},y_{1},\ldots,x_{\lceil g/2\rceil},y_{\lceil g/2\rceil}$ and take a Hamilton path of $G-V_{\delta}$ or $G-V_{\delta}-W$ between $x_{i}$ and $y_{i}$ for each $i\in[1,\lceil g/2\rceil]$, where we will let $w$ be exactly one of those selected vertices if $|W|=1$ and $g_{0}$ is odd. Each pair $x_{i},y_{i}$ will be the two endvertices of a path connecting them. We describe below how to select those vertices.
Let $\alpha^{*}=\Delta/n$. As $|V_{\delta}|<\eta n$, $G-V_{\delta}$ is a robust $(\nu-\eta,2\tau)$-expander by Lemma 2.6. Thus $G-V_{\delta}$ is Hamilton-connected by Lemma 2.7. If $|W|=1$ and $g_{0}$ is odd, we let $x_{1}=w$ and $y_{1}$ be any vertex from $V_{\Delta}$, and let $P_{1}$ be a Hamilton $(x_{1},y_{1})$-path of $G-V_{\delta}$. Otherwise, we let $x_{1},y_{1}\in V_{\Delta}$ be any two distinct vertices, and let $P_{1}$ be a Hamilton $(x_{1},y_{1})$-path of $G-V_{\delta}-W$. Denote $G_{0}=G-V_{\delta}-W-E(P_{1})$.
Suppose $P_{1}=u_{1}u_{2}\ldots u_{p}$, where $p=n-|V_{\delta}|$ if $|W|=1$ and $g_{0}$ is odd and $p=n-|V_{\delta}|-|W|$ otherwise, $u_{1}=x_{1}$ and $u_{p}=y_{1}$. Note that $p\geq g+1$ and $G_{0}$ is an $(\alpha^{*},2\eta)$-almost regular robust $(\nu-2\eta,2\tau)$-expander by Lemma 2.6. For each $i\in[2,\lceil g/2\rceil]$, we let $x_{i}=u_{2(i-1)}$ and $y_{i}=u_{2(i-1)+1}$. Let $\ell=\lceil g/2\rceil-1$.
If $g_{0}\leq 1$, for each $i\in[2,\lceil g/2\rceil]$, we let $F_{i}=x_{i}y_{i}$ be a linear forest. For each $i\in[2,\ell+1]$, let $v_{i1},v_{i2}\in V(G_{0})\setminus V(F_{i})$ be distinct and such that, for any $v\in V(G_{0}^{*})$, there exist at most two $(i,j)\in[2,\ell+1]\times[1,2]$ such that $v=v_{ij}$. For each $i\in[2,\ell+1]$, let
$L_{i}=\{v_{i1}x_{i}y_{i}v_{i2},v_{i2}v_{i1}\}.$
Denote $L=\bigcup_{i\in[2,\ell+1]}L_{i}$. Note that $(L_{2},F_{2}),\dots,(L_{\ell+1},F_{\ell+1})$ are layouts such that, for each $i\in[2,\ell+1]$, $V(L_{i})\subseteq V(G_{0})$, $|V(L_{i})|=2e(F_{i})+2<\eta n$ and $e(L_{i})\leq 2e(F_{i})+3<\eta n$. Moreover, for each $v\in V(G_{0})$, $d_{L}(v)\leq 1+2<\eta n$ and there exist at most $1+2<\eta n$ indices $i\in[2,\ell+1]$ such that $v\in V(L_{i})$.
Applying Theorem 2.12 with $G_{0}$, $\alpha^{*}$, $\nu-2\eta$, $2\tau$, and $\alpha$ playing the roles of $G,\alpha,\nu,\tau$, and $\eta^{*}$, and with $\varepsilon^{*}=2\eta$, $\varepsilon=\eta^{1/5}$, and $\ell=\lceil g/2\rceil-1\leq\frac{1}{2}(\alpha^{*}-\alpha)n$, we find $\ell$ edge-disjoint Hamilton cycles $C_{2},\ldots,C_{\ell+1}$ of $G_{0}+\{x_{i}y_{i}\mid i\in[2,\lceil g/2\rceil]\}$ such that $E(F_{i})\subseteq E(C_{i})$ (as $L_{i}$ has no isolated vertices, a spanning configuration of shape $(L_{i},F_{i})$ is a Hamilton cycle of $G_{0}\cup F_{i}$). Let $P_{i}=C_{i}-x_{i}y_{i}$. Then $P_{i}$ is a Hamilton $(x_{i},y_{i})$-path of $G_{0}-\bigcup_{j=2}^{i-1}E(P_{j})$. By Theorem 2.12, $G_{0}-\bigcup_{i=2}^{\lceil g/2\rceil}E(P_{i})$ is a robust $(\nu^{\prime}/2,8\tau)$-expander.
If $g_{0}\geq 2$, for each $i\in[2,\lceil(g_{0}+1)/2\rceil]$, and for the vertex $w\in W$, we let $w_{i1},w_{i2}$ be two distinct vertices of $V_{\Delta}\setminus\{x_{i},y_{i}\}$ such that they are adjacent in $G$ to $w$, and the edges $ww_{i1},ww_{i2}$ have not been used so far. We let $F_{i}$ be the linear forest consisting of edges $x_{i}y_{i}$ and $w_{i1}w_{i2}$. (This step is to reduce the degree of the vertex $w$ to $\delta$). For each $i\in[\lceil(g_{0}+1)/2\rceil+1,\lceil g/2\rceil]$, we let $F_{i}=x_{i}y_{i}$ be a linear forest.
For each $i\in[2,\ell+1]$, let $v_{i1},v_{i2}\in V(G_{0})\setminus V(F_{i})$ be distinct and such that, for any $v\in V(G_{0})$, there exist at most two $(i,j)\in[2,\ell+1]\times[1,2]$ such that $v=v_{ij}$. For each $i\in[2,\lceil(g_{0}+1)/2\rceil]$, let
$L_{i}=\{v_{i1}x_{i}y_{i}w_{i1}w_{i2}v_{i2},v_{i2}v_{i1}\}.$
For each $i\in[\lceil(g_{0}+1)/2\rceil+1,\lceil g/2\rceil]$, let
$L_{i}=\{v_{i1}x_{i}y_{i}v_{i2},v_{i2}v_{i1}\}.$
Denote $L=\bigcup_{i\in[2,\ell+1]}L_{i}$. Note that $(L_{2},F_{2}),\dots,(L_{\ell+1},F_{\ell+1})$ are layouts such that, for each $i\in[2,\ell+1]$, $V(L_{i})\subseteq V(G_{0})$, $|V(L_{i})|\leq e(F_{i})+4<\eta n$ and $e(L_{i})\leq 2e(F_{i})+3<\eta n$. Moreover, for each $v\in V(G_{0})$, $d_{L}(v)\leq 1+2<\eta n$ and there exist at most $1+2<\eta n$ indices $i\in[1,\ell]$ such that $v\in V(L_{i})$.
Applying Theorem 2.12 with $G_{0}$, $\alpha^{*}$, $\nu-2\eta$, $2\tau$, and $\alpha$ playing the roles of $G,\alpha,\nu,\tau$, and $\eta^{*}$, and with $\varepsilon^{*}=2\eta$, $\varepsilon=\eta^{1/5}$, and $\ell=\lceil g/2\rceil-1\leq\frac{1}{2}(\alpha^{*}-\alpha)n$, we find edge-disjoint Hamilton cycles $C_{2},\ldots,C_{\ell+1}$ of $G_{0}\cup(\bigcup_{i=2}^{\ell+1}F_{i})$ such that $E(F_{i})\subseteq E(C_{i})$. For each $i\in[2,\lceil(g_{0}+1)/2\rceil]$, we let $P_{i}$ be obtained from $C_{i}$ by deleting $x_{i}y_{i}$ and replacing $w_{i1}w_{i2}$ by $w_{i1}ww_{i2}$; and for each $i\in[\lceil(g_{0}+1)/2\rceil+1,\lceil g/2\rceil]$, we let $P_{i}$ be obtained from $C_{i}$ by deleting $x_{i}y_{i}$. Then $P_{i}$ is a Hamilton $(x_{i},y_{i})$-path of $G-V_{\delta}-E(P_{1})-\bigcup_{j=2}^{i-1}E(P_{j})$ if $i\in[2,\lceil(g_{0}+1)/2\rceil]$, and is a Hamilton $(x_{i},y_{i})$-path of $G_{0}-\bigcup_{j=2}^{i-1}E(P_{j})$ otherwise. Furthermore, by Theorem 2.12, $G_{0}-\bigcup_{i=2}^{\lceil g/2\rceil}E(P_{i})$ is a robust $(\nu^{\prime}/2,8\tau)$-expander.
Let $G_{1}=G-\bigcup_{i=1}^{\lceil g/2\rceil}E(P_{i})$ and $Z=\{x_{1},y_{1},\ldots,x_{\lceil g/2\rceil},y_{\lceil g/2\rceil}\}$. Since $|V_{\delta}\cup W|<\eta n+1$ and $G_{0}-\bigcup_{i=2}^{\lceil g/2\rceil}E(P_{i})$ is a robust $(\nu^{\prime}/2,8\tau)$-expander, it follows by Lemma 2.6 that $G_{1}$ is a robust $(\nu^{\prime}/3,16\tau)$-expander.
By the construction above, vertices of $V_{\delta}$ were not involved in the process. Thus we have $d_{G_{1}}(v)=\delta$ for each $v\in V_{\delta}$. Every vertex $v$ of $V_{\Delta}\setminus Z$ was used as an internal vertex for each of the path removed, thus we have $d_{G_{1}}(v)=\Delta-2\lceil g/2\rceil$. If $|W|=0$, then every vertex of $Z$ was used as an endvertex of exactly one of the paths removed, and as an internal vertex of all other paths, thus we have $d_{G_{1}}(v)=\Delta-2\lceil g/2\rceil+1$ for each $v\in Z$. If $|W|=1$, then the vertex $w$ was used exactly $\lceil g_{0}/2\rceil$ times, where exactly $\lceil(g_{0}+1)/2\rceil-1$ times as an internal vertex of $P_{2},\ldots,P_{\lceil(g_{0}+1)/2\rceil}$, and $\lceil g_{0}/2\rceil+1-\lceil(g_{0}+1)/2\rceil$ times as an endvertex of $P_{1}$ ($\lceil g_{0}/2\rceil+1-\lceil(g_{0}+1)/2\rceil=1$ if and only if $g_{0}$ is odd, and is 0 otherwise). As a consequence, we have $d_{G_{1}}(w)=d_{G}(w)-2\lceil g_{0}/2\rceil+\lceil g_{0}/2\rceil+1-\lceil(g_{0}+1)/2\rceil=\delta$.
As edges of $E(P_{1})$ were removed from $G$ when getting $G_{1}$, we know that $E(P_{1})\subseteq E(\overline{G}_{1})$. Also, for $P_{1}$, we have $w\in V(P_{1})$ if and only if $g_{0}$ is odd, and $w=u_{1}$ in this case. Since $Z=\{u_{1},u_{p}\}\cup\{u_{2(i-1)},u_{2(i-1)+1}\mid i\in[2,\lceil g/2\rceil]\}$, we know that $V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})\cap(Z\cup V_{\delta}\cup W)=\emptyset$. Since $u_{2\lceil g/2\rceil}P_{1}u_{p-1}$ is a path, it has a perfect matching if its order is even, and has a matching that covers all but one vertices if its order is odd. Since $|V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})|\geq n-|V_{\delta}|-|Z|-|W|$ and $u_{2\lceil g/2\rceil}P_{1}u_{p-1}\subseteq\overline{G_{1}}$, we then conclude that $\overline{G}_{1}$ has a matching $M_{0}$ that covers at least $n-|V_{\delta}|-|Z|-2$ vertices of $G_{1}$ excluding vertices of $Z\cup V_{\delta}\cup W$. We consider two cases below regarding the parity of $g$ to complete the proof.
Case 3.1: $g$ is even.
Then we have $d_{G_{1}}(v)=\delta+1$ for every $v\in Z$ and $d_{G_{1}}(v)=\delta$ for every $v\in V(G)\setminus Z$. Thus $\Delta(G_{1})=\delta+1$. Since $g$ is even, we know that $\delta+1$ and $\Delta$ have different parities.
If $\Delta$ is odd, then $\delta+1$ is even. As $G_{1}$ has precisely ngn-g vertices of degree $\delta$, it follows that ngn-g is even and so $n$ is even. Also, we know that $n-g\geq\Delta+1-g=\delta+1$. As $n-|V_{\delta}|-|Z|-2\geq n-\delta-g-1$ (since $|V_{\delta}|<\eta n$ and $\delta\geq\alpha n$), $\overline{G}_{1}$ has a matching $M\subseteq M_{0}$ covering exactly $n-g-(\delta+1)$ vertices not including any vertex of $Z$. Let $G_{2}$ be obtained from $G_{1}+M$ by adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup V(M))$. The graph $G_{2}$ is $(\delta+1)$-regular and is a robust $(\nu^{\prime}/3,\tau^{*})$-expander. Thus $G_{2}$ has a Hamilton decomposition by Theorem 1.2. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+1)/2$ linear forests. Thus $\operatorname{la}(G)\leq g/2+(\delta+1)/2=(\Delta+1)/2$, as desired.
If $\Delta$ is even, then $\delta+1$ is odd. If $n$ is odd, then $n-g-(\delta+1)$ is even. We proceed exactly the same way above to obtain a $(\delta+1)$-regular graph $G_{2}$ and obtain $\operatorname{la}(G)\leq g/2+(\delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired. Thus we assume that $n$ is even. As $n-g-\delta-2<n-|V_{\delta}|-|Z|-2$, $\overline{G}_{1}$ has a matching $M_{1}\subseteq M_{0}$ covering exactly $n-g-\delta-2$ vertices of $G_{1}$ avoiding vertices of $Z\cup V_{\delta}\cup W$. We let $G_{2}$ be obtained from $G_{1}+M_{1}$ by adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup V(M_{1}))$. Then we have $d_{G_{2}}(x)=\delta+2$ and $d_{G_{2}}(v)=\delta+1$ for all $v\in V(G_{2})$ with $v\neq x$. Applying Lemma 2.5, $\overline{G_{2}}$ has a matching excluding the vertex $x$ and covering at least $n-\frac{n}{n-\delta-1}-3$ vertices of $G_{2}$. Since $\delta\leq n-3$ by (2) and $g\geq 1$, we have $n-\frac{n}{n-\delta-1}-3\geq n-\delta-2$. Thus $\overline{G_{2}}$ has a matching $M_{2}$ covering exactly $n-\delta-2$ vertices excluding the vertex $x$. Now we let $G_{3}$ be obtained from $G_{2}+M_{2}$ by adding a new vertex $y$ and all the edges joining $y$ and each vertex from $V(G_{2})\setminus(V(M_{2})\cup\{x\})$. The graph $G_{3}$ is $(\delta+2)$-regular and is a robust $(\nu^{\prime}/3,\tau^{*})$-expander. Thus $G_{3}$ has a Hamilton decomposition by Theorem 1.2. Deleting all the edges of $G_{3}$ incident with $x$ or $y$ and all the edges of $M_{1}\cup M_{2}$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+2)/2$ linear forests. Thus $\operatorname{la}(G)\leq g/2+(\delta+2)/2=(\Delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired.
Case 3.2: $g$ is odd.
Then we have $d_{G_{1}}(v)=\delta-1$ for each $v\in V_{\Delta}\setminus Z$, and $d_{G_{1}}(v)=\delta$ for each $v\in V_{\delta}\cup W\cup Z$. Thus $\Delta(G_{1})=\delta$.
If $\delta$ is even, then $\Delta$ is odd. As $G_{1}$ has precisely $n-(g+1)-|V_{\delta}|-|W|$ vertices of odd degrees, which are all $\delta-1$, we know that $n-(g+1)-|V_{\delta}|-|W|$ is even. As $n-|V_{\delta}|-|Z|-2\geq n-(g+1)-\delta-|V_{\delta}|-|W|$, $\overline{G}_{1}$ has a matching $M\subseteq M_{0}$ covering exactly $n-(g+1)-\delta-|V_{\delta}|-|W|$ vertices not including any vertex of $Z\cup W\cup V_{\delta}$. Let $G_{2}$ be obtained from $G_{1}+M$ by adding a new vertex $x$ and all the edges joining $x$ and each vertex from $V(G_{1})\setminus(Z\cup W\cup V_{\delta}\cup V(M))$. The graph $G_{2}$ is $\delta$-regular and is a robust $(\nu^{\prime}/3,\tau^{*})$-expander. Thus $G_{2}$ has a Hamilton decomposition by Theorem 1.2. Deleting all the edges of $G_{2}$ incident with $x$ and all the edges of $M$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $\delta/2$ linear forests. Thus $\operatorname{la}(G)\leq(g+1)/2+\delta/2=(\Delta+1)/2$, as desired.
Next, consider that $\delta$ is odd and so $\Delta$ is even. Then $G_{1}$ has precisely $g+1+|V_{\delta}|+|W|$ vertices of odd degrees, which are all $\delta$, it follows that $g+1+|V_{\delta}|+|W|$ is even. Thus if $n$ is even, then $n-(g+1)-|V_{\delta}|-|W|$ is even. We proceed exactly the same way above to obtain a $\delta$-regular graph $G_{2}$. By Theorem 1.2, we have $\operatorname{la}(G_{2})\leq(\delta+1)/2$. Thus $\operatorname{la}(G)\leq(g+1)/2+(\delta+1)/2=\lceil(\Delta+1)/2\rceil$, as desired.
Thus we assume that $n$ is odd. We claim that $p=n-|V_{\delta}|-|W|$ in this case; that is, we cannot have that $|W|=1$ and $g_{0}$ is odd. This is true, since if $|W|=1$ and $g_{0}$ is odd, then $d_{G}(w)=\delta+g_{0}$ is even. Since $\Delta$ is even, this would imply that $|V_{\delta}|$ must be even and so $|W|$ must be even (since $g+1+|V_{\delta}|+|W|$ is even), a contradiction to $|W|=1$. Then as $g+1+|V_{\delta}|+|W|$ is even, we know that $p=n-|V_{\delta}|-|W|$ is odd. As all the edges in $u_{2\lceil g/2\rceil}P_{1}u_{p-1}$ are contained in $\overline{G}_{1}$, and $V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})=V(G)\setminus(Z\cup V_{\delta}\cup W)$, we let $M_{1}$ be a matching of $\overline{G}_{1}$ covering all the vertices of $V(u_{2\lceil g/2\rceil}P_{1}u_{p-1})\setminus\{u_{p-1}\}$. Let $G_{2}$ be obtained from $G_{1}$ by adding all the edges of $M_{1}$ and the edge $u_{p-1}u_{p}$ from $\overline{G}_{1}$. Then in $G_{2}$, all vertices other than $u_{p}$ have degree $\delta$ and $u_{p}$ has degree $\delta+1$. The graph $\overline{G}_{2}$ has one single vertex of degree $n-1-(\delta+1)$ and all other vertices of degree $n-1-\delta$. By Lemma 2.5, $\overline{G}_{2}$ has a matching covering at least $n-n/(n-\delta)-3$ vertices avoiding the vertex $u_{p}$. As $n$ is odd and $\delta$ is odd, we have $\delta\leq n-2$. As a consequence, $n-n/(n-\delta)-3>n-1-(\delta+1)$. Thus $\overline{G}_{2}$ has a matching $M_{2}$ covering exactly $n-1-(\delta+1)$ vertices excluding the vertex $u_{p}$. We let $G_{3}$ be obtained from $G_{2}+M_{2}$ by adding a new vertex $x$ and all the edges joining $x$ and vertices from $V(G_{2})\setminus(V(M_{2})\cup\{u_{p}\})$. The graph $G_{3}$ is $(\delta+1)$-regular and is a robust $(\nu^{\prime}/3,\tau^{*})$-expander. Thus $G_{3}$ has a Hamilton decomposition by Theorem 1.2. Deleting all the edges of $G_{3}$ incident with $x$ and all the edges of $M_{1}\cup M_{2}$ from the decomposition gives a linear forest decomposition of $G_{1}$ into $(\delta+1)/2$ linear forests. Thus $\operatorname{la}(G)\leq(g+1)/2+(\delta+1)/2=(\Delta+2)/2=\lceil(\Delta+1)/2\rceil$, as desired.
The proof of Theorem 3.1 is now completed. ∎
Acknowledgments
The authors are deeply grateful to the referee, who carefully read the paper, provided us with highly constructive comments, and guided us toward proving a more general result.